### v2

Abstract
We present a high-dynamic-range vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer. This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation. Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ (1 part per billion), as well as two polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength. The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively. Crucially, this approach avoids several metrological difficulties associated with vector magnetometers and gradiometers. We detail the fundamental, systematic, and practical limits of such vector magnetometers. Notably, we provide a comprehensive study of the systematic effects of vector atomic magnetometers. We introduce a new concept of dynamic heading error and investigate several other systematic effects. A unique cosine-altering rotating field modulation is proposed to cancel out these systematics. Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors. This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
keywords:
high dynamic range, vector atomic magnetometry, dynamic heading error, systematics
1 Introduction
Measurements of magnetic fields have a rich history, starting with the invention of the compass. Among the first and still, most widely used vector sensors are fluxgate magnetometers based on asymmetry in saturation of a soft magnetic material in an applied oscillating magnetic field [1]. Advanced condensed-matter-based approaches involve SQUID magnetometers [2], Hall sensors [3] and magneto-resistive sensors [4]. These sensors are intrinsically vector devices that measure one component of the magnetic field, so three separate sensors are typically required for full vector sensing. For example, the ESA Swarm mission necessitates three-axis fluxgate magnetometers and an Absolute Scalar Magnetometer for calibration [5]. In contrast, atomic magnetometers are inherently scalar devices that gauge the energy gap between spin states in a magnetic field. Additional interactions are required to define the vector axes for a scalar sensor.
Sensitive vector magnetometers operating in Earth’s field have several metrological challenges. For example, a 50 fT magnetic field corresponds to 1 part in $10^{9}$ of Earth’s field. Such relative precision is beyond the capability of most analog voltage measurements. However, frequency measurements can easily achieve much higher relative precision. SQUID magnetometers also can achieve higher fractional resolution using fluxquanta counting techniques [2]. Another challenge is a stable alignment and orthogonality of vector axes directions, which require nano-radian stability. One of the advantages of using external magnetic fields is their relative uniformity that allows the same vector axes orientation for multiple sensors. We use an applied rotating field to create vector sensitivity. For any coil imperfections or non-orthogonality, as long as the coils remain linear, a rotating field always has a unique rotation plane that defines the vector axes in our approach.
Atomic vector sensors operating in the spin-exchange relaxation-free (SERF) regime are considered the most sensitive [6]; however, their low dynamic range limits their applications. Scalar magnetometers based on multi-pass cells have also reached the sub-$\mathrm{fT/\sqrt{Hz}}$ sensitivity regime [7]. Scalar atomic magnetometers have been widely used outside the lab in various geomagnetic and space applications [8]. Recently, they have also been used to detect biomagnetic fields outside the lab [9]. Vector sensing of similar sensitivity will provide better localization accuracy for detecting magnetic sources [10], whether in the Earth or the brain.
Several different approaches have created vector sensitivity for atomic sensors. For instance, variometers based on scalar magnetometers [11, 12, 13, 14], magnetic field modulation [15, 16], multi-laser beams to measure multi-spin projections [17, 18, 19, 20], nonlinear magneto-optical rotation (NMOR) magnetometers based on atomic alignment [21, 22], and vector magnetometry based on magneto-optical phenomena [23, 24]. In particular, all-optical vector magnetometers can be achieved by modulation of the light shift, which acts as a pseudo-magnetic field modulation [25]. Lastly, extracting magnetic field vector information by using microwaves as a 3D spatial reference [26]. Some progress in achieving high dynamic range vector magnetometers is listed in Table 1.
Table 1: Vector atomic magnetometers and their performances.
Year Type Tested field ($\mu$T) Total field sensitivity (fT/$\sqrt{\mathrm{Hz}}$) Fractional resolution (1/$\sqrt{\mathrm{Hz}}$) Polar angle sensitivity (nrad/$\sqrt{\mathrm{Hz}}$) Ref
2004 Unshielded SERF magnetometer 50 1,000 (down to 10 Hz) 2×$\times 10^{-8}$ [15]
2006 Variometer based on scalar cesium magnetometer 50 4,000 (estimated) $\times 10^{-8}$ $\times 10^{2}$ (estimated) $\times 10^{2}$ (estimated) [14]
2014 All optical vector NMOR atomic magnetometer 1.5 65 (down to 0.5 Hz) 4×$\times 10^{-8}$ $\times 10^{5}$ only measure one polar angle [25]
2015 Multi-laser beam atomic vector magnetometer based on free spin precession 1 143 (estimated) 1×$\times 10^{-7}$ $4\times 10^{4}$ (estimated) $4\times 10^{4}$ (estimated) [18]
2019 Spin self-sustaining Larmor precession vector magnetometer 0.2 780 (down to 2 Hz) 3.9×$\times 10^{-6}$ 3.5×$\times 10^{3}$ 3.5×$\times 10^{3}$ [27]
2019 Voigt-effect-based vector magnetometer 0.26 400 (down to 2 Hz) 2×$\times 10^{-6}$ 8×$\times 10^{3}$ (down to 10 Hz) 8×$\times 10^{3}$ (down to 10 Hz) [23]
2023 Machine learning assisted NMOR vector magnetometer 0.14 100 (down to 20 Hz) $\times 10^{-7}$ $10^{5}$ (down to 16 Hz) 2×$\times 10^{5}$ (down to 18 Hz) [28]
- FRF vector magnetometer with CARFM modulation 50 50 (down to 1 Hz) 1×$\times 10^{-9}$ 8 (down to 0.1 Hz) 8 (down to 0.1 Hz) This work
We introduce a Fast Rotating Field (FRF) vector magnetometer that can extract three-axis vector information without compromising its scalar resolution. This is achieved by adding a fast rotating magnetic field to compact scalar magnetometers. This approach has enabled the realization of sensitivity as low as several $\mathrm{fT/\sqrt{Hz}}$ in the earth’s ambient environment using multi-pass cells [9, 29].
Our method involves a scalar magnetometer configuration using a pulsed pump laser similar to [9], along with the application of a fast-rotating magnetic field with a rotation rate faster than the spin relaxation rate. This entire sensor can be miniaturized to a very compact size using microelectromechanical systems techniques [30, 31, 32, 33].
We conducted a thorough study on the systematic effects of the magnetometer and developed a unique modulation technique called cosine altering rotating field modulation (CARFM) to cancel out these systematic errors. The magnetometer’s performance is evaluated, and we propose analytical fundamental limits for the sensitivity of the vector free-spin-precession magnetometer based on the Cramer-Rao Lower Bound (CRLB). This is the first comprehensive analysis of the systematic effects of vector atomic magnetometers.
2 Results
2.1 Spin dynamics
The complete description of spin evolution involves density matrix theory [34]. When ground states are distributed according to spin temperature, spin evolution can be effectively described by the Bloch equations [35, 36]. The approximation of rotation matrices is utilized to assess spin evolution in the rotating frame, enabling the derivation of analytical solutions to the time-dependent Bloch equation.
As shown in Fig. 1a, we assume the spins are fully polarized along the negative y-axis. The rotating field initially originates from the x-axis and rotates counterclockwise in the xy-plane (towards the negative z-axis). The spin evolution, without considering spin relaxations, can be expressed as:
$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{\mathbf{e}}_{z}]\cdot\mathcal{R}[\psi,\mathbf{b}]\cdot(-\hat{\mathbf{e}}_{y}),$ (1)
where $\mathcal{R}[\phi,\mathbf{\nu}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{\nu}$. For the applied rotating field in the x-y plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $\text{Abs}[\phi_{x}-\phi_{y}]=\pi/2$. Therefore, $\mathbf{b}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$. Here, $\gamma$ represents the gyromagnetic ratio, and $\hat{\mathbf{e}}_{x}$, $\hat{\mathbf{e}}_{y}$ and $\hat{\mathbf{e}}_{z}$ are the unit vectors along the x, y and z axes, respectively. We define $\theta=\omega_{m}t$ and $\psi=\gamma t\sqrt{B_{z}-\omega_{m}/\gamma}$. Eventually, the spin projections can be written as:
$P_{x}(t)=$ $\hat{\mathbf{e}}_{x}\cdot\mathbf{P}(t)=\cos\omega_{0}t\sin\omega_{m}t+$ (2)
$\frac{\gamma B_{z}-\omega_{m}}{\omega_{0}}\sin\omega_{0}t\cos\omega_{m}t,$
$P_{z}(t)=$ $\hat{\mathbf{e}}_{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (3)
where $\omega_{0}=\gamma\sqrt{B_{m}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq.1, we can get
$P_{x}(t)\approx\cos\left(\omega_{0}t+\gamma\frac{B_{m}b_{y}\cos\omega_{m}t-B_{m}b_{x}\sin\omega_{m}t}{\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}}\right)$ (4)
$\sin\omega_{m}t+\sin\left(\omega_{0}t+\gamma\frac{B_{m}b_{y}\cos\omega_{m}t-B_{m}b_{x}\sin\omega_{m}t}{\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}}\right)$
$\cos\omega_{m}t\cdot\frac{\gamma B_{z}-\omega_{m}}{\omega_{0}},$
the residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift. The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 2 details the sign dependence). Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
b) Four-shot rotating field scheme
a) Experimental setup
c) Phase shift analysis
Figure 1: (a) Three coils determine the x, y, and z directions. The sensor head sits on a rotating stage with the cell positioned at the stage’s center. The sensor head can freely rotate in the x-z and y-z planes. The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$. The angle between the sensor head’s axis and the z-axis is labeled as $\beta$. $\theta$ represents the angle between $\mathbf{B_{tot}}$ and $\mathbf{\hat{e}_{z}}$. Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode. (b) There are four shots with different start phases of $B_{x}$ and $B_{y}$. This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA. The magnetic field only alters during the peak values (“cosine switch”) to suppress the eddy current caused by the rotating field. The signals during the preparation time are blanked out. (c) Upper: When a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in b) are analyzed by the MDA as a function of time. Lower: when the magnetic fields in the x and y directions are roughly compensated. The first harmonic signals are suppressed and the second harmonic signal can be observed.
2.2 Systematics
To achieve high accuracy, systematic errors need to be studied and eliminated. There are several systematics in the FRF vector magnetometer. By generating a power series expansion of Eq. 2 and analyzing the phase shifts, we can identify two phase-shift terms contributed from $B_{m}$. One is Berry’s phase shift, which can be written as
$\tau_{B}(t)=\frac{B_{m}^{2}\omega_{m}t}{2B_{z}^{2}}\approx(1-\cos\theta)\omega_{m}t,$ (5)
where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig. 1a. The Eq. 5 is consistent with Berry’s phase shift described in Ref. [37]. Another phase shift from Eq. 2 is proportional to $\sin(2\omega_{m}t)$, which can be written as
$\tau_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (6)
This phenomenon results in a non-zero second harmonic phase shift even in the absence of a residual magnetic field, and when the rotating field is perfectly symmetric in the x-y plane. Asymmetry in the rotating field along the x and y directions also produces a second harmonic signal.
Regarding the probe beam heading error, we define the angle between the probe beam and the x-axis as $\alpha$, as shown in Fig. 1(b). The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$. By inserting Eqs. 2 and 3, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
$\tau_{prob}(t)\approx-\frac{B_{m}}{B_{z}}\tan\alpha\cdot\sin\omega_{m}t.$ (7)
Converting it into a magnetic field unit, which is equivalent to a fictitious magnetic field $\omega_{m}/\gamma\cdot\tan\alpha$ along the probe beam direction, we named it a probe beam heading error. This error induces a systematic effect where the measured magnetic field depends on the relative angle between the probe beam and $B_{z}$, analogous to the pump beam heading error.
Additionally, there are two extra sources of systematic errors: heading errors and eddy currents. Heading errors include the commonly adopted static heading error, which depends on the sensor’s orientation with respect to the static magnetic field. We introduce the concept of the dynamic heading error, which occurs when the total magnetic field is rotating rather than static. The combined heading error induced by the total magnetic field is the sum of these static and dynamic heading errors. (More details in Sec. 2.3)
Another systematic error is induced by eddy currents. The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field. The magnitude of the eddy current magnetic field highly depends on how the rotating magnetic field is altered. (More details in Sec. 2.3)
The signs of these systematic errors depend on the rotating field’s phases, while their absolute magnitudes remain the same. The sign of the dynamic heading error ($\tau_{DH}$) depends on the rotating field’s phases, but the sign of the static heading error does not. We introduce a four-shot scheme (shown in Fig. 1b) by switching the phases of the rotating field in the order listed in Table 2. By averaging these four shots, the systematic errors can be canceled out.
Table 2: The signs of the systematics as a function of the phases of the rotating magnetic field.
Shot No. $\phi_{x}$ $\phi_{y}$ $b_{x}$ $b_{y}$ $\tau_{B}$ $\tau_{prob}$ $\tau_{DH}$ $\tau_{2nd}$
1 $\pi/2$ 0 - + + - + -
2 $\pi/2$ $\pi$ + + +
3 $23\pi/$ 2 $\pi$ + - + - - -
4 $23\pi/$ 2 $2\pi$ + + - + - +
2.3 Experimental results
To study systematics, a single-pass alkali vapor cell is placed at the center of a magnetic shield, which is a cubic cell with internal dimensions of 5$\times$5$\times$5 mm3. The cell contains a droplet of 87Rb and 688 Torr N2 as quenching and buffer gas, respectively. The magnetic shield comprises a two-layer mu-metal shield and an innermost-layer aluminum shield designed to attenuate DC and AC magnetic fields from the environment.
The alkali cell is pumped by a sequence of pump pulses from a grating-stabilized diode laser. Inside the magnetic shield, a set of three coils provides a rotating magnetic field and a leading magnetic field. The cell is heated by an electric heater driven by AC at a frequency of 131.5 kHz. The AC heater is turned on during the pump time and turned off during the measurement time to reduce magnetic noise from the heater.
As shown in Fig. 1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction. The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$. The frequency of the pump beam pulse is set to 348 kHz, synchronized with the Larmor frequency to achieve the highest pumping rate. The duty cycle of a pump beam pulse is approximately 1.4%.
After the pump time, the pump beam is turned off, and the free induction decay (FID) signal of the spin precession is measured by a linearly polarized probe beam originating from a vertical cavity surface emitting laser (VCSEL). The optical rotation of the linearly polarized probe beam is measured by a balanced polarimeter consisting of a Wollaston prism and a quadrant photodiode. The signal from the photodiode is then passed through a differential low-noise amplifier and a high-pass filter with a cutoff frequency of 150 kHz.
The phase shift of the acquired signal relative to the Larmor frequency is analyzed using the HP 53310A modulation domain analyzer (MDA), which detects the zero-crossings of the precession signal and a reference signal of $\omega_{\text{ref}}$ that is close to the precession frequency. The phase shifts between the zero crossings of the precession signal and the reference signal are then calculated. The phase shift measuremment of one block (four shots) of data is shown in Fig. 1c.
The fitting model further fits the measured phase shifts $\delta\Phi(t)$ from the MDA as a function of measurement time to get: offset, slope ($\Phi_{s}$, Table A1 in the supplemental material details the components of the slope), amplitudes of the first harmonics ($b_{x}\propto\Phi_{x}$ and $b_{y}\propto\Phi_{y}$), second harmonics of $\omega_{m}$, and amplitudes of the first harmonics of the hyperfine frequency $\omega_{\text{hp}}$ (the difference in precession frequency for the F=1 and F=2 atoms induces an exponentially decaying sinusoidal phase shift at a frequency of $\omega_{\text{hp}}=2\pi\times 1.39$ kHz, see supplemental material Sec. I).
$\begin{split}\delta\Phi(t)=&\Phi_{off}+\Phi_{s}\cdot t+\Phi_{x}\sin{\omega_{m}t}+\Phi_{y}\cos{\omega_{m}t}\\ &+\Phi_{2s}\sin{2\omega_{m}t}+\Phi_{2c}\cos{2\omega_{m}t}\\ &+\Phi_{hps}\sin{\omega_{hp}t}+\Phi_{hpc}\cos{\omega_{hp}t}\end{split}$ (8)
[…]
2.3.1 Results of the systematics
[Probe beam heading error] To study the probe beam heading error, the sensor rotates about the y-axis in the x-z plane with an angle of $\alpha$. This error can induce a magnetic field measurement systematic that depends on the angle $\alpha$, and its sign depends on the rotating direction of the rotating field. The measured probe heading error when $\omega_{m}=2\pi\times 480$ Hz is shown in Fig. 2(a). The theory is based on Eq. 7. “Theory 1” is for anti-clockwise rotations (Shot 1 and Shot 3 in Table 2), which is $-\omega_{m}/\gamma\cdot\tan\alpha$. “Theory 2” is for clockwise rotations (Shots 2 and 4 in Table 2), which is $\omega_{m}/\gamma\cdot\tan\alpha$. The probe beam heading error can be canceled out by averaging two opposite rotating field measurements.
[Berry’s phase] When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 5 predicts 4.7 nT, as shown in Table. 3.
[Second harmonic] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 6 prediction well.
[Static heading error] The static heading error is measured when the magnetometer sensor rotates an angle $\beta$ = 24 in the y-z plane which leads to a heading error $B_{SH}=\mathcal{B}_{H}\sin\beta$, we set nuclear Landé factor $g_{I}\approx 0$ [38] and
$\mathcal{B}_{H}\approx\frac{P(7+P^{2})}{5+3P^{2}}\frac{3h\gamma B_{tot}^{2}}{4\pi A_{hf}},$ (9)
where $P$ is spin polarization, $A_{hf}=h\cdot 3.417$ GHz is the hyperfine structure constant for ground state [39], $h$ is Planck constant. We can estimate a static heading error of 2.7 nT based on these equations (P=78%, the initial polarization is estimated using the decay feature of the FID signal envelope [40]), the experimental result is 2.7 nT. The static heading error is independent of the rotating field.
Table 3: The measurement results and the predictions are based on the rotating matrix simulation and density matrix simulation (static heading error).
Systematic Measured Predicted
Berry’s phase 5.1 nT 4.7 nT
Amplitude of $\sin 2\omega_{m}t$ 35 mrad 35 mrad
Static heading error $B_{SH}$ 2.7 nT 2.7 nT
(a) Probe beam heading error
(b) Dynamic heading error
(c) Eddy current due to different switches
(d) Eddy current time constant
Figure 2: (a) The measurement of the probe beam heading error. The solid points are experimental results, and the dashed lines are the theoretic plots based on our proposed model. (b) The magnetic field in the transverse direction is caused by the eddy current. The magnetic fields measured by both the Sine and Cosine switches are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches. (d) The measurement of the eddy current time constant.
[Dynamic heading error] In addition to the static heading error, a rotating magnetic field can induce a dynamic heading error. This effect is characterized in Table 4 using experimental and density matrix simulation results [34]. The density matrix model describes an ensemble of spins in a mixed state and is detailed in Secs. VI and VII of the supplemental material.
When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be explained because the spin precession plane can adiabatically follow the rotation of the total magnetic field $\mathbf{B_{tot}}$. This keeps the relative angle $\Theta$ (as depicted in Fig. 2(b)) between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a constant precession frequency due to the weighted sum of different Zeeman transitions stays consistent. Thus, the dynamic heading error $B_{DH}=\mathcal{B}_{H}\cos\Theta$ depends on the initial angle $\Theta$ between the spin polarization and $\mathbf{B_{tot}}$, which remains constant and equal to its initial value.
Immediately after laser pumping, the initial spin $\mathbf{S}$ is along the negative y direction. $\mathbf{S_{\parallel}}$ and $\mathbf{S_{\perp}}$ are the spin projections of $\mathbf{S}$ along and perpendicular to the total magnetic field $\mathbf{B_{tot}}$ immediately after optical pumping, respectively.
If the $\mathbf{B_{tot}}$ starts initially perpendicular to $\mathbf{S}$ ( $\mathbf{S_{\parallel}}=0$), such as when $\mathbf{B_{m}}$ starts from the x-axis and $\mathbf{B_{tot}}$ is in the x-z plane, $\mathbf{S}$ is initially perpendicular to $\mathbf{B_{tot}}$ ($\Theta=\pi/2$), resulting in a dynamic heading error of 0. The spin precession plane can adiabatically follow the rotation of the total magnetic field, remaining perpendicular to it.
However, if the $\mathbf{B_{tot}}$ initially starts non-perpendicularly to $\mathbf{S}$ ($\mathbf{S_{\parallel}}\neq 0$), such as when $\mathbf{B_{m}}$ starts from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the y-z plane, $\mathbf{S}$ is no longer perpendicular to $\mathbf{B_{tot}}$, and there is an angle of $\Theta=\pi/2-\theta\approx 69^{\circ}$. This angle leads to a measured dynamic heading error of 2.6 nT, which is close to $\mathcal{B}_{H}\cos\Theta$. Similarly, if the rotating field starts from the positive y-axis ($\Theta=\pi/2+\theta$), the dynamic heading error becomes -2.6 nT. The full sign dependence of the dynamic heading error can be found in Table 2.
Table 4: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$. “Exp” is the experimental data, “DM” is the density matrix simulation result.
$\mathbf{B_{m}}$ starts $\parallel\mathbf{S}$ $\mathbf{B_{m}}$ starts $\perp\mathbf{S}$
$B_{tot}$ Shot No. Exp. (nT) DM (nT) Exp. (nT) DM (nT)
1 2.6 3.0 0.1 0.0
2 2.6 3.0 0.4 0.0
3 -2.6 -3.0 -0.1 0.0
4 -2.6 -3.0 -0.4 0.0
Avg 0.0 0.0 0.0 0.0
$B_{x}$ 1 -1.5 -0.1 -0.8 0.0
2 0.4 0.1 0.8 0.0
3 -0.3 -0.1 -1.0 0.0
4 1.6 0.0 1.3 0.0
Avg 0.0 0.0 0.1 0.0
$B_{y}$ 1 -0.1 0.1 -0.1 0.0
2 0.5 0.1 0.1 0.0
3 0.1 -0.1 0.1 0.0
4 -0.2 -0.1 0.2 0.0
Avg 0.1 0.1 0.1 0.0
The experimental results for the dynamic heading error are smaller than the density matrix simulation because the actual spin polarization is approximately 22% smaller than the simulation, which assumes near 100% polarization. For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec. VII of the supplemental material.
We also evaluated the dynamic heading error effect in the transverse directions. The summary of the experimental results and the theoretical predictions related to $\mathbf{B_{tot}}$ based on the density matrix are shown in Table 4. The dynamic heading error effect is close to zero in $B_{y}$ regardless of whether the rotating field starts parallel or perpendicular to $\mathbf{S}$. These small magnetic field variations may be caused by systematic mismatches in the threshold voltage of the MDA.
The variation in $B_{x}$ measurement from the density matrix simulation in each shot is attributed to the probe beam heading error ($\tau_{\text{prob}}$). For instance, if the sensor has a misalignment such that $\alpha=1^{\circ}$, it leads to an approximate 1.2 nT probe beam heading error in the x-direction. However, these systematic errors can be eliminated through 4-shot averaging.
Additional data in Sec. VII of the supplemental material show that these magnetic field offsets do not change their signs when the polarization direction is reversed. This proves that they are not caused by the dynamic heading error.
Overall, the rotating magnetic field can induce dynamic heading errors in total magnetic field measurements. These errors can be eliminated by either choosing the right starting phases for the rotating field, which starts the field initially perpendicular to the spin polarization, or by averaging the results of four shots, as shown in Table 2. The dynamic heading error does not contribute to the transverse magnetic fields, regardless of the starting phases of the rotating field.
[Eddy current] The fast-altering rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield. In our four-shot scheme that alternates the rotation direction, there are two switches in four shots, as shown in Fig. 1b. We discovered that how the rotating magnetic field is altered can affect the eddy current and its induced magnetic field. When the direction alteration of the rotating field happens at the peak values of the rotating magnetic field, from negative maximum to positive maximum and vice versa, we named this as the “cosine switch” (Fig. 1b top left). The average magnetic field caused by the eddy current equals zero. However, if the direction alteration of the rotating field happens at zero rotating magnetic field, we name it the “sine switch” (Fig. 1b top right), which would lead to a significant magnetic field caused by the eddy current.
For instance, in y-direction, in Fig. 1b, the switch of $B_{y}$ occurs during the preparation time of the second shot, and another switch occurs during the preparation time of the fourth shot. Therefore, in the “sine switch” case, the second and fourth shots have the maximum eddy current, while the first and third shots have smaller eddy currents because they measure the decaying eddy current generated during the fourth and second shots. The sign of the magnetic field caused by the eddy current depends on the rotation direction of the rotating field. One panorama (80 shots) measurement result from the MDA is shown in Fig. 2(c). The sine switch can cause an eddy current magnetic field of approximately 450 nT, while the cosine switch can highly suppress the systematic of the eddy current. The small eddy current magnetic field in the cosine switch is due to the imperfection of the cosine switch. The instruments have a certain rising time for a sudden signal transition.
The time constant of the eddy current is measured by applying a continuous rotating magnetic field with a constant phase (no switch) during one panorama measurement. As shown in Fig. 2(d), there are 79 data points representing 79 shot measurements. The first shot is dumped due to it overloading the phase shift range of the MDA. For each shot, it takes 1/120 s. The 60 Hz line frequency causes periodic variations in the magnified picture and its harmonics, which can be averaged out by one panorama measurement. We measured the time constant of the eddy current effect for our setup to be approximately 10.4 ms, with the corresponding cut-off frequency at 15 Hz.
Therefore, we specifically developed a waveform as shown in Fig. 1b to ensure the rotation direction alteration always happens at the peak values of the adjacent rotating magnetic field (cosine switch). We can mostly eliminate the systematic effect caused by the eddy current by using the ”cosine switch” and 4-shot averaging.
2.3.2 Vector magnetic gradiometer
To demonstrate the magnetic gradiometer measurement of the FRF vector magnetometer, a multipass 87Rb cell is used. A high-power QPC laser polarizes the spins along the z-axis to increase spin the polarization, and $\pi/2$ magnetic field pulses are generated right after the pump to tip the spins from the z-axis to the y-axis. Two Carmel NK732 cards replace the MDA to acquire the rising and falling edge zero-crossings for 1 minute. The gradiometer sensitivity is shown in Fig. 3, with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ for measuring the total field and sensitivities of 400 $\mathrm{fT/\sqrt{Hz}}$ for measuring $B_{x}$ and $B_{y}$, respectively. Their noise spectrums are flat down to 1 Hz and 0.1 Hz, respectively. For comparison, a scalar sensitivity of 40 $\mathrm{fT/\sqrt{Hz}}$ was achieved without applying any rotating field.
Figure 3: Magnetic gradiometer sensitivity. The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and two transverse magnetic fields. The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer. The inset displays the frequency response of $B_{x}$ and $B_{y}$.
3 Discussion
In summary, we introduce a FRF vector geomagnetic magnetometer. This system is achieved by applying a rotating field to a scalar atomic magnetometer, enabling it to fully determine the magnetic field vector. While the modulation slightly degrades the scalar performance from 40 $\mathrm{fT/\sqrt{Hz}}$ to 50 $\mathrm{fT/\sqrt{Hz}}$, it provides two additional polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$. This enhancement allows the vector magnetometer to precisely measure two transverse polar angles, eliminating the static heading error common in scalar atomic magnetometers. The unique vector axes are defined by the rotation plane of the applied field, increasing stability compared to relying on mechanical coil orthogonality.
We provide a comprehensive study of the systematics and develop cosine-altering rotating field modulation to cancel out these effects. Benefits from frequency measurement, such vector magnetometers can achieve high fractional resolution. Additionally, the vector magnetometer retains the accuracy and metrological advantages of scalar atomic magnetometers, such as inherent calibration.
Furthermore, the fundamental sensitivity of the vector magnetometer in total field measurement is identical to that of scalar magnetometers. Compared with approaches using sequential modulation, our proposed vector magnetometer is faster (higher bandwidth) and experiences fewer systematic effects. The FRF vector magnetometer can be further improved to reach quantum-limited sensitivity [41].
4 Methods
4.1 Noise limit
Based on the CRLB [42] and the photon shot noise of the balanced polarimeter detection [43], the power spectral density $\rho(\omega)$ has a minimum when the measurement time $t\approx 2T_{2}$ (check Sec. III of supplemental material for more details). The total magnetic field measurement sensitivity is given by
$\delta B_{tot}\geq\frac{4}{\gamma kT_{2}\sqrt{\Phi_{pr}}}.$ (10)
where $k=lr_{e}cfnD(\nu)/2$, $l$ is the length of probe beam travel through the cell, $r_{e}$ is the classical electron radius, $f$ is the typical oscillator strength of the D line transition, $D(\nu)=(\nu-\nu_{0})/[(\nu-\nu_{0})^{2}+(\Delta\nu/2)^{2}]$, where $\Delta\nu$ is the optical full width at half maximum (FWHM), $\nu_{0}$ is the frequency of the D line transition. $T_{2}$ is the transverse relaxation time. $\Phi_{pr}$ is the total flux of photons of the probe beam per unit of time. If we keep the measurement time the same as the optimized measurement time of the total magnetic field $t=2T_{2}$, then the transverse magnetic field measurement sensitivity is given by
$\delta B_{tran}\geq\frac{2\sqrt{2}\omega_{m}\csc\theta}{\gamma k\sqrt{\Phi_{pr}}}.$ (11)
The derived fundamental sensitivities for measuring the scalar magnetic field and transverse magnetic field are based on the assumption that the fitting of these fields is independent. However, this assumption is not accurate when the modulation frequency is not high enough. An analytical solution for the sensitivities can be obtained by fitting the scalar and transverse magnetic fields simultaneously. This can be achieved using the Minimum-Variance Unbiased (MVU) estimator, similar to the approach described in Sections III and IV of the supplemental material. The results are more complex, but it is found that when the modulation frequency is fast enough, the assumption of independent fitting holds. ($\omega_{m}\gtrsim\pi/T_{2}$), the fundamental sensitives are almost the same as considering them independently.
For comparison, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers can be described as (Please refer to Sec. IV of the supplemental material for details.)
$\delta B_{tran\_SM}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (12)
If we assume $\omega_{m}=\pi/T_{2}$, the sensitivity of the transverse magnetic field of the FRF vector magnetometer can be written as
$\delta B_{tran\_FRF}=\frac{2\sqrt{2}\pi}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (13)
The FRF vector magnetometer demonstrates a sensitivity in measuring transverse magnetic fields that is comparable to sequential modulation vector magnetometers. Additionally, it offers a higher bandwidth and is less susceptible to systematic effects.
Acknowledgements
The authors thank Dmitry Budker, Conlon Lorcan Oneill for the valuable comments. The authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA), A*STAR Career Development Fund ( [^2] D800028), Italy-Singapore science and technology collaboration grant (R23I0IR042), Delta-Q (C230917004, Quantum Sensing), and Competitive Research Programme (NRF-CRP30-2023-0002).
Author contributions statement
T.W. prepared the setup and conducted the experiments for the FRF vector magnetometers, proposed and implemented the CARFM modulation, modeled the spin evolution using rotation matrices and density matrices, derived the fundamental limits, analyzed the results, and prepared the manuscript. M.R. supervised the setup and experiments, modeled the spin evolution using rotation matrices, and discussed the CARFM modulation and dynamic heading error. W.L. assisted with modeling the static heading error and analyzing the polarization of the atoms. M.L., T.K., and E.F. provided the scalar magnetometer sensors and offered some valuable experimental advice. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.
Additional information
The corresponding author is responsible for submitting a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.

### v3

@@
-
+Pulsed Vector Atomic Magnetometer Using an Alternating Fast-Rotating Field
 Abstract
-We present a high-dynamic-range vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer.
+We present a vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer.
 This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation.
-Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ (1 part per billion), as well as two polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength.
+Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 part per billion), as well as two polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength.
 The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively.
@@
 We introduce a new concept of dynamic heading error and investigate several other systematic effects.
-A unique cosine-altering rotating field modulation is proposed to cancel out these systematics.
+A unique peak-altering fast rotating field modulation is proposed to cancel out these systematics.
 Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors.
 This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
-keywords:
 high dynamic range, vector atomic magnetometry, dynamic heading error, systematics
-1 Introduction
-Measurements of magnetic fields have a rich history, starting with the invention of the compass.
+preprint: APS/123-QEP
+I Introduction
+The measurement of magnetic fields has a long and rich history, beginning with the invention of the compass.
+The earliest known compass, the “south-governor” (sinan), dates back over 2000 years.
 Among the first and still, most widely used vector sensors are fluxgate magnetometers based on asymmetry in saturation of a soft magnetic material in an applied oscillating magnetic field [1].
@@
 For example, the ESA Swarm mission necessitates three-axis fluxgate magnetometers and an Absolute Scalar Magnetometer for calibration [5].
-In contrast, atomic magnetometers are inherently scalar devices that gauge the energy gap between spin states in a magnetic field.
+In contrast, atomic magnetometers are inherently scalar devices that gauge the energy gap between spin states in a magnetic field, detecting a frequency directly proportional to the field strength.
+This frequency-based measurement approach allows scalar magnetometers to achieve very high relative precision.
 Additional interactions are required to define the vector axes for a scalar sensor.
@@
 Another challenge is a stable alignment and orthogonality of vector axes directions, which require nano-radian stability.
-One of the advantages of using external magnetic fields is their relative uniformity that allows the same vector axes orientation for multiple sensors.
 We use an applied rotating field to create vector sensitivity.
@@
 Lastly, extracting magnetic field vector information by using microwaves as a 3D spatial reference [26].
-Some progress in achieving high dynamic range vector magnetometers is listed in Table 1.
-Table 1: Vector atomic magnetometers and their performances.
-Year Type Tested field ($\mu$T) Total field sensitivity (fT/$\sqrt{\mathrm{Hz}}$) Fractional resolution (1/$\sqrt{\mathrm{Hz}}$) Polar angle sensitivity (nrad/$\sqrt{\mathrm{Hz}}$) Ref
-2004 Unshielded SERF magnetometer 50 1,000 (down to 10 Hz) 2×$\times 10^{-8}$ [15]
-2006 Variometer based on scalar cesium magnetometer 50 4,000 (estimated) $\times 10^{-8}$ $\times 10^{2}$ (estimated) $\times 10^{2}$ (estimated) [14]
-2014 All optical vector NMOR atomic magnetometer 1.5 65 (down to 0.5 Hz) 4×$\times 10^{-8}$ $\times 10^{5}$ only measure one polar angle [25]
-2015 Multi-laser beam atomic vector magnetometer based on free spin precession 1 143 (estimated) 1×$\times 10^{-7}$ $4\times 10^{4}$ (estimated) $4\times 10^{4}$ (estimated) [18]
-2019 Spin self-sustaining Larmor precession vector magnetometer 0.2 780 (down to 2 Hz) 3.9×$\times 10^{-6}$ 3.5×$\times 10^{3}$ 3.5×$\times 10^{3}$ [27]
-2019 Voigt-effect-based vector magnetometer 0.26 400 (down to 2 Hz) 2×$\times 10^{-6}$ 8×$\times 10^{3}$ (down to 10 Hz) 8×$\times 10^{3}$ (down to 10 Hz) [23]
-2023 Machine learning assisted NMOR vector magnetometer 0.14 100 (down to 20 Hz) $\times 10^{-7}$ $10^{5}$ (down to 16 Hz) 2×$\times 10^{5}$ (down to 18 Hz) [28]
-- FRF vector magnetometer with CARFM modulation 50 50 (down to 1 Hz) 1×$\times 10^{-9}$ 8 (down to 0.1 Hz) 8 (down to 0.1 Hz) This work
+While all of these methods measure the amplitudes or frequencies of Larmor precession, frequency measurements can achieve higher fractional precision and are less affected by drifts in cell temperature, pump and probe laser powers, and wavelengths.
+We focus on measuring the zero-crossings of the Larmor precession signals, which not only provide frequency information but also capture phase shifts and integrate magnetic field modulations, offering more comprehensive insights into the magnetic fields.
 We introduce a Fast Rotating Field (FRF) vector magnetometer that can extract three-axis vector information without compromising its scalar resolution.
-This is achieved by adding a fast rotating magnetic field to compact scalar magnetometers.
-This approach has enabled the realization of sensitivity as low as several $\mathrm{fT/\sqrt{Hz}}$ in the earth’s ambient environment using multi-pass cells [9, 29].
-Our method involves a scalar magnetometer configuration using a pulsed pump laser similar to [9], along with the application of a fast-rotating magnetic field with a rotation rate faster than the spin relaxation rate.
+This is achieved by adding a fast rotating magnetic field—at a rotation rate exceeding the spin relaxation rate—to compact scalar magnetometers.
+Utilizing multipass cells, these scalar magnetometers achieve sensitivities as low as several $\mathrm{fT/\sqrt{Hz}}$ in the earth’s ambient environment [9, 27].
+Our method involves a scalar magnetometer configuration using a pulsed pump laser similar to [9], along with the application of a fast-rotating magnetic field.
 This entire sensor can be miniaturized to a very compact size using microelectromechanical systems techniques [28, 29, 30, 31].
-We conducted a thorough study on the systematic effects of the magnetometer and developed a unique modulation technique called cosine altering rotating field modulation (CARFM) to cancel out these systematic errors.
+We conducted a thorough study on the systematic effects of the magnetometer and developed a unique modulation technique called peak-altering fast rotating field modulation to cancel out these systematic errors.
 The magnetometer’s performance is evaluated, and we propose analytical fundamental limits for the sensitivity of the vector free-spin-precession magnetometer based on the Cramer-Rao Lower Bound (CRLB).
-This is the first comprehensive analysis of the systematic effects of vector atomic magnetometers.
-2 Results
-2.1 Spin dynamics
+II Results
+II.1 Spin dynamics
 The complete description of spin evolution involves density matrix theory [32].
 When ground states are distributed according to spin temperature, spin evolution can be effectively described by the Bloch equations [33, 34].
+$\frac{d\mathbf{P}}{dt}=\gamma\mathbf{P}\times\mathbf{B},$ (1)
+where $\gamma$ represents the gyromagnetic ratio, $\mathbf{P}$ is spin polarization, and $\mathbf{B}$ is magnetic field.
 The approximation of rotation matrices is utilized to assess spin evolution in the rotating frame, enabling the derivation of analytical solutions to the time-dependent Bloch equation.
@@
 1a, we assume the spins are fully polarized along the negative y-axis.
-The rotating field initially originates from the x-axis and rotates counterclockwise in the xy-plane (towards the negative z-axis).
+The magnetic field initially points along the z-axis, while a rotating field originates from the x-axis and rotates counterclockwise in the xy-plane.
+We consider the behavior of polarization in a magnetic field, where it undergoes precession in response to the applied field.
+This magnetic field is composed of both a static and a rotating component.
+In the presence of these components, the polarization vector evolves according to specific rotations.
 The spin evolution, without considering spin relaxations, can be expressed as:
-$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{\mathbf{e}}_{z}]\cdot\mathcal{R}[\psi,\mathbf{b}]\cdot(-\hat{\mathbf{e}}_{y}),$ (1)
-where $\mathcal{R}[\phi,\mathbf{\nu}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{\nu}$.
-For the applied rotating field in the x-y plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $\text{Abs}[\phi_{x}-\phi_{y}]=\pi/2$.
-Therefore, $\mathbf{b}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$.
-Here, $\gamma$ represents the gyromagnetic ratio, and $\hat{\mathbf{e}}_{x}$, $\hat{\mathbf{e}}_{y}$ and $\hat{\mathbf{e}}_{z}$ are the unit vectors along the x, y and z axes, respectively.
-We define $\theta=\omega_{m}t$ and $\psi=\gamma t\sqrt{B_{z}-\omega_{m}/\gamma}$.
+$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{B_{tot}}]\cdot(-\hat{y}),$ (2)
+where $\mathcal{R}[\phi,\mathbf{v}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{v}$.
+For the applied rotating field in the xy-plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$.
+Therefore, $\mathbf{B_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$.
+Here, $\hat{x}$, $\hat{y}$ and $\hat{z}$ are the unit vectors along the x, y and z axes, respectively.
+We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{B_{tot}}|$.
 Eventually, the spin projections can be written as:
-$P_{x}(t)=$ $\hat{\mathbf{e}}_{x}\cdot\mathbf{P}(t)=\cos\omega_{0}t\sin\omega_{m}t+$ (2)
+$P_{x}(t)=$ $\hat{x}\cdot\mathbf{P}(t)=\cos\omega_{0}t\sin\omega_{m}t+$ (3)
 $\frac{\gamma B_{z}-\omega_{m}}{\omega_{0}}\sin\omega_{0}t\cos\omega_{m}t,$
-$P_{z}(t)=$ $\hat{\mathbf{e}}_{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (3)
-where $\omega_{0}=\gamma\sqrt{B_{m}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq.1, we can get
+$P_{z}(t)=$ $\hat{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (4)
+where $\omega_{0}=\gamma|\mathbf{B_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
 $P_{x}(t)\approx\cos\left(\omega_{0}t+\gamma\frac{B_{m}b_{y}\cos\omega_{m}t-B_{m}b_{x}\sin\omega_{m}t}{\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}}\right)$ (5)
@@
 $\cos\omega_{m}t\cdot\frac{\gamma B_{z}-\omega_{m}}{\omega_{0}},$
+where $|\mathbf{B_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$.
 The residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift.
@@
 Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
-b) Four-shot rotating field scheme
+The intuitive classical model underlying the operating principle of the FRF vector magnetometer is illustrated in Fig.
+1a.
+The scalar magnitude of the total magnetic field, $|\mathbf{B_{tot}}|$, depends on the angle between the rotating field $\mathbf{B_{m}}$ and the residual magnetic field $\mathbf{B_{0}}$.
+When a small residual transverse magnetic field $\mathbf{\Delta B}$—comprising $\mathbf{b_{x}}$ and $\mathbf{b_{y}}$—is present, $|\mathbf{B_{tot}}|$ varies with oscillating components at the rotation frequency of $\mathbf{B_{m}}$.
+The phases and amplitudes of these oscillations are influenced by the angle between $\mathbf{B_{0}}$ and $\mathbf{B_{m}}$.
+Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an in-phase oscillation, while a residual field in the y-direction induces an out-of-phase component.
+This model provides a fundamental description of the magnetometer’s operation.
+However, for precise measurements, several systematic effects—some even resulting from hyperfine interactions—require in-depth analysis to account for and minimize potential measurement biases.
 a) Experimental setup
 c) Phase shift analysis
+b) Four-shot rotating field scheme
 Figure 1: (a) Three coils determine the x, y, and z directions.
 The sensor head sits on a rotating stage with the cell positioned at the stage’s center.
-The sensor head can freely rotate in the x-z and y-z planes.
+The sensor head can freely rotate in the xz and yz-planes.
 The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$.
-The angle between the sensor head’s axis and the z-axis is labeled as $\beta$.
-$\theta$ represents the angle between $\mathbf{B_{tot}}$ and $\mathbf{\hat{e}_{z}}$.
+$\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field.
+$\theta$ represents angle between the axis of rotation and the rotating field vector.
 Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode.
-(b) There are four shots with different start phases of $B_{x}$ and $B_{y}$.
+(b) Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms.
+When low after a falling edge, the pump laser and AC heater are turned off, and the probe beam laser measures the FID signal.
+There are four shots with different start phases of $B_{x}$ and $B_{y}$.The start phases of the rotating field play a significant role; therefore, we indicate them with vertical dashed lines.
 This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA.
-The magnetic field only alters during the peak values (“cosine switch”) to suppress the eddy current caused by the rotating field.
+The magnetic field only alters during the peak values (“Peak-Altering”) to suppress the eddy current caused by the rotating field.
 The signals during the preparation time are blanked out.
@@
 The first harmonic signals are suppressed and the second harmonic signal can be observed.
-2.2 Systematics
-To achieve high accuracy, systematic errors need to be studied and eliminated.
-There are several systematics in the FRF vector magnetometer.
-By generating a power series expansion of Eq.
-2 and analyzing the phase shifts, we can identify two phase-shift terms contributed from $B_{m}$.
-One is Berry’s phase shift, which can be written as
-$\tau_{B}(t)=\frac{B_{m}^{2}\omega_{m}t}{2B_{z}^{2}}\approx(1-\cos\theta)\omega_{m}t,$ (5)
+II.2 Systematics
+To achieve high accuracy, systematic errors must be carefully studied and eliminated.
+In the FRF vector magnetometer, several systematic effects are present.
+By expanding Eq.
+3 into a power series and examining the phase shifts to second order in $B_{m}$, we can identify two phase-shift contributions from $B_{m}$:
+[Berry’s phase shift] The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35].
+As the magnetic field rotates slowly, the spins adiabatically follow the field.
+At the end of a complete cycle, when the field returns to its original configuration, the spins acquire a phase shift.
+If the system undergoes cyclic adiabatic evolution at a constant frequency, the Berry phase accumulates linearly over time.
+Thus, the Berry phase term as a function of time, from Eq.
+3, can be written as
+$\phi_{B}(t)=\frac{B_{m}^{2}\omega_{m}t}{2B_{z}^{2}}\approx(1-\cos\theta)\omega_{m}t,$ (6)
 where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig.
 1a.
-The Eq.
-5 is consistent with Berry’s phase shift described in Ref.
+The Berry’s phase acquired over one complete period, given by Eq.
+6 as $2\pi(1-\cos{\theta})$, is consistent with Berry’s phase shift of Nuclear Magnetic Resonance (NMR) described in Ref.
 [35].
-Another phase shift from Eq.
-2 is proportional to $\sin(2\omega_{m}t)$, which can be written as
-$\tau_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (6)
-This phenomenon results in a non-zero second harmonic phase shift even in the absence of a residual magnetic field, and when the rotating field is perfectly symmetric in the x-y plane.
-Asymmetry in the rotating field along the x and y directions also produces a second harmonic signal.
-Regarding the probe beam heading error, we define the angle between the probe beam and the x-axis as $\alpha$, as shown in Fig.
+[Second harmonic phase shift] Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
+$\phi_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (7)
+This phenomenon leads to a non-zero second harmonic phase shift, even in the absence of a residual magnetic field and when the rotating field is perfectly symmetric in the xy-plane.
+Any asymmetry in the rotating field along the x and y-directions also produces a second harmonic signal instead of the first harmonic at $\omega_{m}$, ensuring that it does not affect the measurement of transverse magnetic fields.
+[Heading errors] The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36].
+The pump beam heading error is well-studied and is primarily caused by nonlinear Zeeman splitting and the difference between Zeeman resonances in the two hyperfine ground states.
+We set nuclear Landé factor $g_{I}\approx 0$ [37] and the heading error can be well described by $B_{SH}=\mathcal{B_{H}}\sin\beta$, where
+$\mathcal{B_{H}}\approx\frac{P(7+P^{2})}{5+3P^{2}}\frac{3h\gamma B_{tot}^{2}}{4\pi A_{hf}},$ (8)
+$\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field.
+$A_{hf}=h\cdot 3.417$ GHz is the hyperfine structure constant for ground state [38], $h$ is Planck constant.
+We refer it to as the static heading error (to distinguish it from the dynamic heading error introduced below, we refer to this conventional heading error as static heading error), which depends on the sensor’s orientation with respect to the static magnetic field.
+We introduce the concept of the dynamic heading error, which occurs when the total magnetic field is rotating rather than static.
+We discovered that the dynamic heading error can be explained by the fact that the spin precession plane adiabatically follows the rotation of the total magnetic field.
+This maintains a constant relative angle between the total magnetic field vector and the spin precession plane, resulting in a stable precession frequency, as the weighted sum of different Zeeman transitions remains consistent.
+Consequently, the dynamic heading error depends on the initial angle between the spin polarization and the total magnetic field, which remains constant and equal to its initial value.
+(More details in Sec.
+II.3).
+To clarify the elimination of the systematic dynamic heading error, we separate the dynamic heading error, $B_{DH}$, into two components: a static part ($B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{PD}$).
+Notably, the static heading error is independent of the rotating field phases.
+There’s an analogous “heading error” effect from the probe beam.
+This error creates a systematic effect where the measured transverse magnetic field depends on the relative angle between the probe beam and the plane of the rotating field, analogous to the pump beam heading error.
+Regarding the probe beam heading error, we define the angle that the probe beam rotates away from the x-axis as $\alpha$, with the sensor rotating in the xz-plane., as shown in Fig.
 1(b).
@@
 By inserting Eqs.
-2 and 3, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
-$\tau_{prob}(t)\approx-\frac{B_{m}}{B_{z}}\tan\alpha\cdot\sin\omega_{m}t.$ (7)
-Converting it into a magnetic field unit, which is equivalent to a fictitious magnetic field $\omega_{m}/\gamma\cdot\tan\alpha$ along the probe beam direction, we named it a probe beam heading error.
-This error induces a systematic effect where the measured magnetic field depends on the relative angle between the probe beam and $B_{z}$, analogous to the pump beam heading error.
-Additionally, there are two extra sources of systematic errors: heading errors and eddy currents.
-Heading errors include the commonly adopted static heading error, which depends on the sensor’s orientation with respect to the static magnetic field.
-We introduce the concept of the dynamic heading error, which occurs when the total magnetic field is rotating rather than static.
-The combined heading error induced by the total magnetic field is the sum of these static and dynamic heading errors.
-(More details in Sec.
-2.3)
-Another systematic error is induced by eddy currents.
+3 and 4, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
+$\phi_{prob}(t)\approx-\frac{B_{m}}{B_{z}}\tan\alpha\cdot\sin\omega_{m}t.$ (9)
+Compared to Eq.
+5, this is equivalent to transverse magnetic field $b_{x}$.
+Converting it into a magnetic field unit, it corresponds to a fictitious magnetic field $-\omega_{m}/\gamma\cdot\tan\alpha$ along the x-direction, which we refer to as a probe beam heading error.
+[Eddy current] Another systematic error is induced by eddy currents.
 The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field.
@@
 (More details in Sec.
-2.3)
-The signs of these systematic errors depend on the rotating field’s phases, while their absolute magnitudes remain the same.
-The sign of the dynamic heading error ($\tau_{DH}$) depends on the rotating field’s phases, but the sign of the static heading error does not.
+II.3)
+[Sign dependence of systematics] We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged.
+This characteristic offers an opportunity to eliminate these systematics.
+The sign of the $B_{PD}$ part of dynamic heading error depends on the rotating field’s phases, but the sign of the static heading error does not.
 We introduce a four-shot scheme (shown in Fig.
-1b) by switching the phases of the rotating field in the order listed in Table 2.
+1b) by altering the phases of the rotating field in the order listed in Table 1.
 By averaging these four shots, the systematic errors can be canceled out.
 Table 1: The signs of the systematics as a function of the phases of the rotating magnetic field.
-Shot No. $\phi_{x}$ $\phi_{y}$ $b_{x}$ $b_{y}$ $\tau_{B}$ $\tau_{prob}$ $\tau_{DH}$ $\tau_{2nd}$
+Shot No. $\phi_{x}$ $\phi_{y}$ $b_{x}$ $b_{y}$ $\phi_{B}$ $\phi_{prob}$ $B_{PD}$ $\phi_{2nd}$
 1 $\pi/2$ 0 - + + - + -
@@
 4 $23\pi/$ 2 $2\pi$ + + - + - +
-2.3 Experimental results
+II.3 Experimental results
 To study systematics, a single-pass alkali vapor cell is placed at the center of a magnetic shield, which is a cubic cell with internal dimensions of 5$\times$5$\times$5 mm3.
@@
 The magnetic shield comprises a two-layer mu-metal shield and an innermost-layer aluminum shield designed to attenuate DC and AC magnetic fields from the environment.
-The alkali cell is pumped by a sequence of pump pulses from a grating-stabilized diode laser.
+The alkali cell is pumped by a sequence of pulses from a grating-stabilized diode laser, with a transverse relaxation time measured at $T_{2}\approx 3$ ms.
 Inside the magnetic shield, a set of three coils provides a rotating magnetic field and a leading magnetic field.
-The cell is heated by an electric heater driven by AC at a frequency of 131.5 kHz.
+The cell is heated to 100C by an electric heater driven by AC at a frequency of 131.5 kHz.
 The AC heater is turned on during the pump time and turned off during the measurement time to reduce magnetic noise from the heater.
 As shown in Fig.
-1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction.
+1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction.
 The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$.
@@
 The duty cycle of a pump beam pulse is approximately 1.4%.
-After the pump time, the pump beam is turned off, and the free induction decay (FID) signal of the spin precession is measured by a linearly polarized probe beam originating from a vertical cavity surface emitting laser (VCSEL).
+After the pump phase, the pump beam is switched off, and the free induction decay (FID) signal of the spin precession is measured using a linearly polarized probe beam.
+The probe beam, with a cross-sectional area of approximately 4 mm2 and a power of around 2 mW, is generated by a vertical cavity surface-emitting laser (VCSEL).
 The optical rotation of the linearly polarized probe beam is measured by a balanced polarimeter consisting of a Wollaston prism and a quadrant photodiode.
@@
 […]
-2.3.1 Results of the systematics
-[Probe beam heading error] To study the probe beam heading error, the sensor rotates about the y-axis in the x-z plane with an angle of $\alpha$.
-This error can induce a magnetic field measurement systematic that depends on the angle $\alpha$, and its sign depends on the rotating direction of the rotating field.
-The measured probe heading error when $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.
-2(a).
-The theory is based on Eq.
-7.
-“Theory 1” is for anti-clockwise rotations (Shot 1 and Shot 3 in Table 2), which is $-\omega_{m}/\gamma\cdot\tan\alpha$.
-“Theory 2” is for clockwise rotations (Shots 2 and 4 in Table 2), which is $\omega_{m}/\gamma\cdot\tan\alpha$.
-The probe beam heading error can be canceled out by averaging two opposite rotating field measurements.
-[Berry’s phase] When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 5 predicts 4.7 nT, as shown in Table. 3.
-[Second harmonic] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 6 prediction well.
-[Static heading error] The static heading error is measured when the magnetometer sensor rotates an angle $\beta$ = 24 in the y-z plane which leads to a heading error $B_{SH}=\mathcal{B}_{H}\sin\beta$, we set nuclear Landé factor $g_{I}\approx 0$ [38] and
-$\mathcal{B}_{H}\approx\frac{P(7+P^{2})}{5+3P^{2}}\frac{3h\gamma B_{tot}^{2}}{4\pi A_{hf}},$ (9)
-where $P$ is spin polarization, $A_{hf}=h\cdot 3.417$ GHz is the hyperfine structure constant for ground state [39], $h$ is Planck constant.
-We can estimate a static heading error of 2.7 nT based on these equations (P=78%, the initial polarization is estimated using the decay feature of the FID signal envelope [40]), the experimental result is 2.7 nT.
-The static heading error is independent of the rotating field.
-Table 3: The measurement results and the predictions are based on the rotating matrix simulation and density matrix simulation (static heading error).
+II.3.1 Results of the systematics
+The experimental results for the systematics are presented in Table 2.
+Table 2: The measurement results and predictions are based on rotating matrix simulations and analytical models.
 Systematic Measured Predicted
@@
 Amplitude of $\sin 2\omega_{m}t$ 35 mrad 35 mrad
-Static heading error $B_{SH}$ 2.7 nT 2.7 nT
+Static heading error $B_{SH}$ ($\beta\approx 24^{\circ}$) 2.7 nT 2.9 nT
+$|B_{PD}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$) 2.6 nT 2.8 nT
+[Berry’s phase] The Berry phase will result in a phase shift that adds to the total magnetic field measurement.
+When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq.
+6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table.
+2.
+[Second harmonic phase shift] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
+[Heading Errors] The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$.
+From Eq.
+(8), the static heading error is estimated to be 2.9 nT, with an experimental result of 2.7 nT.
+The slightly lower experimental value is attributed to incomplete spin polarization.
 (a) Probe beam heading error
@@
 (b) The magnetic field in the transverse direction is caused by the eddy current.
-The magnetic fields measured by both the Sine and Cosine switches are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches.
+The magnetic fields measured by both the Zero-Altering and Peak-Altering are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches.
 (d) The measurement of the eddy current time constant.
-[Dynamic heading error] In addition to the static heading error, a rotating magnetic field can induce a dynamic heading error.
-This effect is characterized in Table 4 using experimental and density matrix simulation results [34].
-The density matrix model describes an ensemble of spins in a mixed state and is detailed in Secs.
-VI and VII of the supplemental material.
-When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be explained because the spin precession plane can adiabatically follow the rotation of the total magnetic field $\mathbf{B_{tot}}$.
-This keeps the relative angle $\Theta$ (as depicted in Fig.
-2(b)) between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a constant precession frequency due to the weighted sum of different Zeeman transitions stays consistent.
-Thus, the dynamic heading error $B_{DH}=\mathcal{B}_{H}\cos\Theta$ depends on the initial angle $\Theta$ between the spin polarization and $\mathbf{B_{tot}}$, which remains constant and equal to its initial value.
-Immediately after laser pumping, the initial spin $\mathbf{S}$ is along the negative y direction.
-$\mathbf{S_{\parallel}}$ and $\mathbf{S_{\perp}}$ are the spin projections of $\mathbf{S}$ along and perpendicular to the total magnetic field $\mathbf{B_{tot}}$ immediately after optical pumping, respectively.
-If the $\mathbf{B_{tot}}$ starts initially perpendicular to $\mathbf{S}$ ( $\mathbf{S_{\parallel}}=0$), such as when $\mathbf{B_{m}}$ starts from the x-axis and $\mathbf{B_{tot}}$ is in the x-z plane, $\mathbf{S}$ is initially perpendicular to $\mathbf{B_{tot}}$ ($\Theta=\pi/2$), resulting in a dynamic heading error of 0.
-The spin precession plane can adiabatically follow the rotation of the total magnetic field, remaining perpendicular to it.
-However, if the $\mathbf{B_{tot}}$ initially starts non-perpendicularly to $\mathbf{S}$ ($\mathbf{S_{\parallel}}\neq 0$), such as when $\mathbf{B_{m}}$ starts from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the y-z plane, $\mathbf{S}$ is no longer perpendicular to $\mathbf{B_{tot}}$, and there is an angle of $\Theta=\pi/2-\theta\approx 69^{\circ}$.
-This angle leads to a measured dynamic heading error of 2.6 nT, which is close to $\mathcal{B}_{H}\cos\Theta$.
-Similarly, if the rotating field starts from the positive y-axis ($\Theta=\pi/2+\theta$), the dynamic heading error becomes -2.6 nT.
-The full sign dependence of the dynamic heading error can be found in Table 2.
+A rotating magnetic field can induce a rotating-field-phase-dependent heading error, $B_{PD}$.
+This effect is characterized in Table 3 with both experimental results and density matrix simulations [32].
+The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Secs.
+VIII and IX of the supplemental material.
+When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$.
+This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency.
+Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value.
+We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig.
+2(b).
+For example, when $\mathbf{B_{m}}$ begins along the x-axis, the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane.
+Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same.
+In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{PD}=0$.
+Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant.
+Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{DH}=0$.
+However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$.
+This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta\approx 2.9$ nT, and $B_{DH}=\mathcal{B_{H}}\cos\beta\sin\theta\approx 2.5~{}\mathrm{nT}$.
+Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{SH}$ remaining the same, but $B_{DH}=-\mathcal{B_{H}}\cos\beta\sin\theta\approx-2.5~{}\mathrm{nT}$.
+Table 1 provides a full overview of the sign dependence of the dynamic heading error.
 Table 3: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$.
+Here the static part of dynamic heading error $B_{SH}=0$.
 “Exp” is the experimental data, “DM” is the density matrix simulation result.
-$\mathbf{B_{m}}$ starts $\parallel\mathbf{S}$ $\mathbf{B_{m}}$ starts $\perp\mathbf{S}$
+$\mathbf{B_{m}}$ initiates from Y-axis $\mathbf{B_{m}}$ initiates from X-axis ($\mathbf{B_{tot}}\perp\mathbf{S}$)
 $B_{tot}$ Shot No. Exp. (nT) DM (nT) Exp. (nT) DM (nT)
@@
 Avg 0.1 0.1 0.1 0.0
-The experimental results for the dynamic heading error are smaller than the density matrix simulation because the actual spin polarization is approximately 22% smaller than the simulation, which assumes near 100% polarization.
+The experimental results for the dynamic heading error are smaller than the analytical model and density matrix simulation because the actual spin polarization is smaller than the simulation, which assumes full polarization.
 For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec.
-VII of the supplemental material.
-We also evaluated the dynamic heading error effect in the transverse directions.
-The summary of the experimental results and the theoretical predictions related to $\mathbf{B_{tot}}$ based on the density matrix are shown in Table 4.
-The dynamic heading error effect is close to zero in $B_{y}$ regardless of whether the rotating field starts parallel or perpendicular to $\mathbf{S}$.
-These small magnetic field variations may be caused by systematic mismatches in the threshold voltage of the MDA.
-The variation in $B_{x}$ measurement from the density matrix simulation in each shot is attributed to the probe beam heading error ($\tau_{\text{prob}}$).
-For instance, if the sensor has a misalignment such that $\alpha=1^{\circ}$, it leads to an approximate 1.2 nT probe beam heading error in the x-direction.
-However, these systematic errors can be eliminated through 4-shot averaging.
+IX of the supplemental material.
+We further evaluated the dynamic heading error effect in the transverse directions.
+A summary of the experimental results and the theoretical predictions related to $\mathbf{B_{tot}}$, based on the density matrix, is provided in Table 3.
+The dynamic heading error effect remains close to zero in $B_{y}$, irrespective of whether the rotating field begins parallel or perpendicular to $\mathbf{S}$.
+These minor magnetic field variations may result from systematic mismatches in the MDA’s threshold voltage, leading to curvature that affects the fitting of the slope $\Phi_{s}$.
+The variation in $B_{x}$ measurement from the density matrix simulation in each shot is due to the probe beam heading error ($\phi_{\text{prob}}/\gamma$).
+For example, if the sensor is misaligned with an angle of $\alpha=1^{\circ}$, this results in an approximate 1.2 nT probe beam heading error in the x-direction.
 Additional data in Sec.
-VII of the supplemental material show that these magnetic field offsets do not change their signs when the polarization direction is reversed.
-This proves that they are not caused by the dynamic heading error.
-Overall, the rotating magnetic field can induce dynamic heading errors in total magnetic field measurements.
-These errors can be eliminated by either choosing the right starting phases for the rotating field, which starts the field initially perpendicular to the spin polarization, or by averaging the results of four shots, as shown in Table 2.
-The dynamic heading error does not contribute to the transverse magnetic fields, regardless of the starting phases of the rotating field.
-[Eddy current] The fast-altering rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield.
-In our four-shot scheme that alternates the rotation direction, there are two switches in four shots, as shown in Fig.
-1b.
-We discovered that how the rotating magnetic field is altered can affect the eddy current and its induced magnetic field.
-When the direction alteration of the rotating field happens at the peak values of the rotating magnetic field, from negative maximum to positive maximum and vice versa, we named this as the “cosine switch” (Fig.
-1b top left).
-The average magnetic field caused by the eddy current equals zero.
-However, if the direction alteration of the rotating field happens at zero rotating magnetic field, we name it the “sine switch” (Fig.
-1b top right), which would lead to a significant magnetic field caused by the eddy current.
-For instance, in y-direction, in Fig.
-1b, the switch of $B_{y}$ occurs during the preparation time of the second shot, and another switch occurs during the preparation time of the fourth shot.
-Therefore, in the “sine switch” case, the second and fourth shots have the maximum eddy current, while the first and third shots have smaller eddy currents because they measure the decaying eddy current generated during the fourth and second shots.
-The sign of the magnetic field caused by the eddy current depends on the rotation direction of the rotating field.
-One panorama (80 shots) measurement result from the MDA is shown in Fig.
-2(c).
-The sine switch can cause an eddy current magnetic field of approximately 450 nT, while the cosine switch can highly suppress the systematic of the eddy current.
-The small eddy current magnetic field in the cosine switch is due to the imperfection of the cosine switch.
-The instruments have a certain rising time for a sudden signal transition.
-The time constant of the eddy current is measured by applying a continuous rotating magnetic field with a constant phase (no switch) during one panorama measurement.
-As shown in Fig.
-2(d), there are 79 data points representing 79 shot measurements.
-The first shot is dumped due to it overloading the phase shift range of the MDA.
-For each shot, it takes 1/120 s.
-The 60 Hz line frequency causes periodic variations in the magnified picture and its harmonics, which can be averaged out by one panorama measurement.
-We measured the time constant of the eddy current effect for our setup to be approximately 10.4 ms, with the corresponding cut-off frequency at 15 Hz.
-Therefore, we specifically developed a waveform as shown in Fig.
-1b to ensure the rotation direction alteration always happens at the peak values of the adjacent rotating magnetic field (cosine switch).
-We can mostly eliminate the systematic effect caused by the eddy current by using the ”cosine switch” and 4-shot averaging.
-2.3.2 Vector magnetic gradiometer
-To demonstrate the magnetic gradiometer measurement of the FRF vector magnetometer, a multipass 87Rb cell is used.
-A high-power QPC laser polarizes the spins along the z-axis to increase spin the polarization, and $\pi/2$ magnetic field pulses are generated right after the pump to tip the spins from the z-axis to the y-axis.
-Two Carmel NK732 cards replace the MDA to acquire the rising and falling edge zero-crossings for 1 minute.
-The gradiometer sensitivity is shown in Fig.
-3, with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ for measuring the total field and sensitivities of 400 $\mathrm{fT/\sqrt{Hz}}$ for measuring $B_{x}$ and $B_{y}$, respectively.
+IX of the supplemental material indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error.
+These systematic errors, however, can be eliminated through 4-shot averaging.
+Overall, the rotating magnetic field can induce dynamic heading errors in total magnetic field measurements, consisting of a static heading error component and a rotating-field-phase-dependent component.
+The rotating-field-phase-dependent heading error, $B_{PD}$, can be eliminated either by selecting appropriate starting phases for the rotating field—ensuring the field initially aligns perpendicularly or maintains a constant angle relative to the spins with alternating start directions—or by averaging the results of four shots, as shown in Table 1.
+Importantly, the dynamic heading error does not affect measurement of the transverse magnetic fields, regardless of the starting phases of the rotating field.
+To investigate the probe beam heading error, the sensor rotates about the y-axis within the xz-plane at an angle of $\alpha$.
+This error can introduce a systematic effect in transverse magnetic field measurements that depends on the angle $\alpha$, with its sign determined by the rotation direction of the rotating field.
+The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.2(a), with the theoretical basis provided in Eq.9.
+“Theory 1” corresponds to anti-clockwise rotations (Shots 1 and 3 in Table 1), given by $-\omega_{m}/\gamma\cdot\tan\alpha$, while “Theory 2” represents clockwise rotations (Shots 2 and 4 in Table 1), given by $\omega_{m}/\gamma\cdot\tan\alpha$.
+The probe beam heading error can be effectively canceled by averaging the measurements from two opposite rotating field directions.
+[Eddy current] The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield.
+In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig.1b.
+We found that the way the rotating magnetic field is altered affects the eddy current and its induced magnetic field.
+When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig.1b, top left).
+In this case, the average magnetic field caused by the eddy current equals zero.
+However, if the direction change occurs when the rotating magnetic field is at zero, termed “Zero-Altering” (Fig.
+1b, top right), it leads to a significant magnetic field caused by the eddy current.
+In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig.1b.
+In the “Zero-Altering” case, the second and fourth shots experience the maximum eddy current, while the first and third shots have smaller eddy currents, as they capture the decaying eddy current generated during the fourth and second shots.
+The sign of the eddy current-induced magnetic field depends on the rotation direction of the rotating field.
+A panorama measurement (80 shots) result from the MDA, shown in Fig.2(c), indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT.
+In contrast, Peak-Altering effectively suppresses the systematic effect of the eddy current.
+Any remaining small eddy current magnetic field in Peak-Altering is due to imperfections, as instruments have a finite rise time during abrupt signal transitions.
+To characterize the time constant of the eddy current, we applied a continuous rotating magnetic field with a constant phase (no alteration) over one panorama measurement.
+In Fig.
+2(d), 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA.
+Each shot measurement lasts 1/120 s.
+The 60 Hz line frequency produces periodic variations and harmonics, which are effectively averaged out over one panorama.
+Our setup’s measured eddy current time constant is approximately 10.4 ms, corresponding to a cut-off frequency of 15 Hz.
+To mitigate the eddy current’s systematic effect, we designed a waveform shown in Fig.
+1b to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering).
+By employing Peak-Altering and 4-shot averaging, we effectively reduce the eddy current’s systematic effect.
+II.3.2 Results of magnetic field sensitivities
+A magnetic field gradiometer measures the variation in magnetic field strength over a short distance, rather than the absolute field strength at a single location.
+This method is effective for detecting local magnetic changes while reducing the influence of uniform, distant magnetic interference [39].To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output.
+The laser, operating in pulsed mode, polarizes the spins along the z-axis to enhance spin alignment and is then turned off during the readout phase.
+Following this pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis.
+Data acquisition is performed with two Carmel NK732 cards in place of the MDA, recording the zero-crossing data—including both frequency and phase information—from two magnetometers over a one-minute interval.
+We derive the gradiometer signal by subtracting the zero-crossing data from the two magnetometers, forming a first-order gradiometer that cancels common-mode magnetic field noise.
+Assuming that any remaining noise is uncorrelated, we further divide this residual noise by $\sqrt{2}$ to assess the intrinsic magnetic field sensitivity of each channel [39].
+The gradiometer sensitivities are shown in Fig.
+3, with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ for the total field and 280 $\mathrm{fT/\sqrt{Hz}}$ for the $B_{x}$ and $B_{y}$ components, respectively.
 Their noise spectrums are flat down to 1 Hz and 0.1 Hz, respectively.
-For comparison, a scalar sensitivity of 40 $\mathrm{fT/\sqrt{Hz}}$ was achieved without applying any rotating field.
-Figure 3: Magnetic gradiometer sensitivity.
-The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and two transverse magnetic fields.
+For comparison, a scalar sensitivity of 28 $\mathrm{fT/\sqrt{Hz}}$ was achieved without the application of a rotating field.
+Figure 3: Magnetic gradiometer sensitivity: The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields.
 The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer.
-The inset displays the frequency response of $B_{x}$ and $B_{y}$.
-3 Discussion
-In summary, we introduce a FRF vector geomagnetic magnetometer.
+The inset shows the frequency response for $B_{x}$ and $B_{y}$.
+III Discussion
+In summary, we present a FRF vector geomagnetic magnetometer.
 This system is achieved by applying a rotating field to a scalar atomic magnetometer, enabling it to fully determine the magnetic field vector.
-While the modulation slightly degrades the scalar performance from 40 $\mathrm{fT/\sqrt{Hz}}$ to 50 $\mathrm{fT/\sqrt{Hz}}$, it provides two additional polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$.
+While the modulation slightly degrades the scalar performance from 28 $\mathrm{fT/\sqrt{Hz}}$ to 35 $\mathrm{fT/\sqrt{Hz}}$, it provides two additional polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$.
 This enhancement allows the vector magnetometer to precisely measure two transverse polar angles, eliminating the static heading error common in scalar atomic magnetometers.
 The unique vector axes are defined by the rotation plane of the applied field, increasing stability compared to relying on mechanical coil orthogonality.
-We provide a comprehensive study of the systematics and develop cosine-altering rotating field modulation to cancel out these effects.
+We provide a comprehensive study of the systematics and develop peak-altering fast rotating field modulation to cancel out these effects.
 Benefits from frequency measurement, such vector magnetometers can achieve high fractional resolution.
@@
 The FRF vector magnetometer can be further improved to reach quantum-limited sensitivity [40].
-4 Methods
-4.1 Noise limit
-Based on the CRLB [42] and the photon shot noise of the balanced polarimeter detection [43], the power spectral density $\rho(\omega)$ has a minimum when the measurement time $t\approx 2T_{2}$ (check Sec.
-III of supplemental material for more details).
-The total magnetic field measurement sensitivity is given by
-$\delta B_{tot}\geq\frac{4}{\gamma kT_{2}\sqrt{\Phi_{pr}}}.$ (10)
-where $k=lr_{e}cfnD(\nu)/2$, $l$ is the length of probe beam travel through the cell, $r_{e}$ is the classical electron radius, $f$ is the typical oscillator strength of the D line transition, $D(\nu)=(\nu-\nu_{0})/[(\nu-\nu_{0})^{2}+(\Delta\nu/2)^{2}]$, where $\Delta\nu$ is the optical full width at half maximum (FWHM), $\nu_{0}$ is the frequency of the D line transition.
-$T_{2}$ is the transverse relaxation time.
-$\Phi_{pr}$ is the total flux of photons of the probe beam per unit of time.
-If we keep the measurement time the same as the optimized measurement time of the total magnetic field $t=2T_{2}$, then the transverse magnetic field measurement sensitivity is given by
+IV Methods
+IV.1 Shot noise limit
+To derive the shot noise limit for the FRF magnetometer, we utilize the CRLB [41] along with photon shot noise considerations specific to balanced polarimeter detection [42].
+Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Sec.
+IV of the supplemental material).
+For total magnetic field measurements, the sensitivity is expressed as
+$\delta B_{tot}\geq\frac{4}{\gamma kT_{2}\sqrt{\Phi_{pr}}},$ (11)
+where $k=lr_{e}cfnD(\nu)/2$, with $l$ being the probe beam path length through the cell, $r_{e}$ the classical electron radius, $f$ the typical oscillator strength of the D-line transition, and $D(\nu)=(\nu-\nu_{0})/[(\nu-\nu_{0})^{2}+(\Delta\nu/2)^{2}]$, where $\Delta\nu$ is the optical full width at half maximum (FWHM) and $\nu_{0}$ is the D-line transition frequency.
+Here, $T_{2}$ is the transverse relaxation time, and $\Phi_{pr}$ denotes the photon flux per unit time.
+This noise limit also applies to scalar atomic magnetometers when analyzing the free induction decay (FID) signal.
+If the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
 $\delta B_{tran}\geq\frac{2\sqrt{2}\omega_{m}\csc\theta}{\gamma k\sqrt{\Phi_{pr}}}.$ (12)
-The derived fundamental sensitivities for measuring the scalar magnetic field and transverse magnetic field are based on the assumption that the fitting of these fields is independent.
-However, this assumption is not accurate when the modulation frequency is not high enough.
-An analytical solution for the sensitivities can be obtained by fitting the scalar and transverse magnetic fields simultaneously.
-This can be achieved using the Minimum-Variance Unbiased (MVU) estimator, similar to the approach described in Sections III and IV of the supplemental material.
-The results are more complex, but it is found that when the modulation frequency is fast enough, the assumption of independent fitting holds.
-($\omega_{m}\gtrsim\pi/T_{2}$), the fundamental sensitives are almost the same as considering them independently.
-For comparison, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers can be described as (Please refer to Sec.
-IV of the supplemental material for details.)
-$\delta B_{tran\_SM}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (12)
-If we assume $\omega_{m}=\pi/T_{2}$, the sensitivity of the transverse magnetic field of the FRF vector magnetometer can be written as
-$\delta B_{tran\_FRF}=\frac{2\sqrt{2}\pi}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (13)
-The FRF vector magnetometer demonstrates a sensitivity in measuring transverse magnetic fields that is comparable to sequential modulation vector magnetometers.
-Additionally, it offers a higher bandwidth and is less susceptible to systematic effects.
+These derived sensitivities assume independent fitting of scalar and transverse magnetic fields.
+However, this assumption becomes inaccurate if the modulation frequency is not sufficiently high.
+Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Sections III and IV of the supplemental material.
+The results show that when modulation frequency is sufficiently fast ($\omega_{m}\gtrsim\pi/T_{2}$), independent fitting yields comparable sensitivities.
+Based on experimental parameters, we estimate the shot noise sensitivities of the FRF magnetometer to be $\delta B_{tot}=0.04~{}\mathrm{fT/\sqrt{Hz}}$ and $\delta B_{tran}=0.8~{}\mathrm{fT/\sqrt{Hz}}$.
+However, experimental measurements show sensitivities worse than these theoretical limits, primarily due to magnetic noise from current sources.
+For instance, an SRS DC205 supplies approximately 650 mA to generate the primary magnetic field along the z-axis, simulating a 50 $\mu$T Earth field.
+Despite modifications to reduce current noise, it remains a primary noise source, limiting overall sensitivity.
+To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Sec. VI of the supplemental material for details):
+$\delta B_{tran\_SM}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}.$ (13)
+If $\omega_{m}=\pi/T_{2}$, the sensitivity for transverse magnetic fields in the FRF vector magnetometer can be expressed as
+$\delta B_{tran\_FRF}=\frac{2\sqrt{2}\pi}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}.$ (14)
+Thus, the FRF vector magnetometer achieves transverse field sensitivity comparable to that of sequential modulation vector magnetometers while offering a higher bandwidth and reduced susceptibility to systematic effects.
 Acknowledgements
@@
 The authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA), A*STAR Career Development Fund ( [^2] D800028), Italy-Singapore science and technology collaboration grant (R23I0IR042), Delta-Q (C230917004, Quantum Sensing), and Competitive Research Programme (NRF-CRP30-2023-0002).
-Author contributions statement
+Author Contributions
 T.W.
-prepared the setup and conducted the experiments for the FRF vector magnetometers, proposed and implemented the CARFM modulation, modeled the spin evolution using rotation matrices and density matrices, derived the fundamental limits, analyzed the results, and prepared the manuscript.
+prepared the setup and conducted experiments for the FRF vector magnetometers, proposed and implemented the modulation, modeled spin evolution using rotation and density matrices, derived fundamental limits, analyzed results, and prepared the manuscript.
 M.R.
-supervised the setup and experiments, modeled the spin evolution using rotation matrices, and discussed the CARFM modulation and dynamic heading error.
+supervised the setup and experiments, contributed to spin evolution modeling with rotation matrices, and discussed modulation and dynamic heading error.
 W.L.
-assisted with modeling the static heading error and analyzing the polarization of the atoms.
+assisted with modeling static heading error and analyzing atomic polarization.
 M.L., T.K., and E.F.
-provided the scalar magnetometer sensors and offered some valuable experimental advice.
+provided scalar magnetometer sensors, offered valuable experimental advice, and provided constructive comments for manuscript revisions.
 All authors reviewed the manuscript.
@@
 The authors declare no competing interests.
-Additional information
-The corresponding author is responsible for submitting a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.

### v4

@@
 Abstract
-We present a vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer.
-This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation.
-Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 part per billion), as well as two polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength.
-The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively.
-Crucially, this approach avoids several metrological difficulties associated with vector magnetometers and gradiometers.
-We detail the fundamental, systematic, and practical limits of such vector magnetometers.
-Notably, we provide a comprehensive study of the systematic effects of vector atomic magnetometers.
-We introduce a new concept of dynamic heading error and investigate several other systematic effects.
-A unique peak-altering fast rotating field modulation is proposed to cancel out these systematics.
-Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors.
-This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
-high dynamic range, vector atomic magnetometry, dynamic heading error, systematics
-preprint: APS/123-QEP
-I Introduction
+We introduce a vector atomic magnetometer that employs a fast-rotating magnetic field applied to a pulsed 87Rb scalar atomic magnetometer.
+This approach enables simultaneous measurements of the total magnetic field and its two polar angles relative to the rotation plane.
+Operating in gradiometer mode, the magnetometer achieves a total field gradient sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 parts per billion) and angular resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at a 50 $\mu$T Earth field strength.
+The noise spectra remain flat down to 1 Hz and 0.1 Hz, respectively.
+Here we show that this method overcomes several metrological challenges commonly faced by vector magnetometers and gradiometers.
+We propose a unique peak-altering modulation technique to mitigate systematic effects, including a newly identified dynamic heading error.
+Additionally, we establish the fundamental sensitivity limits of the sensor, demonstrating that its vector sensitivity approaches scalar sensitivity while preserving the inherent accuracy and calibration benefits of scalar sensors.
+This high-dynamic-range, ultrahigh-resolution magnetometer offers exceptional versatility for diverse applications.
+1 Introduction
 The measurement of magnetic fields has a long and rich history, beginning with the invention of the compass.
@@
 The magnetometer’s performance is evaluated, and we propose analytical fundamental limits for the sensitivity of the vector free-spin-precession magnetometer based on the Cramer-Rao Lower Bound (CRLB).
-II Results
-II.1 Spin dynamics
+2 Results
+2.1 Spin dynamics
 The complete description of spin evolution involves density matrix theory [32].
@@
 As shown in Fig.
-1a, we assume the spins are fully polarized along the negative y-axis.
+1, we assume the spins are fully polarized along the negative y-axis.
 The magnetic field initially points along the z-axis, while a rotating field originates from the x-axis and rotates counterclockwise in the xy-plane.
+We have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$.
+For this assumption, $\phi_{x}=\pi/2$ and $\phi_{y}=0$.
 We consider the behavior of polarization in a magnetic field, where it undergoes precession in response to the applied field.
@@
 The spin evolution, without considering spin relaxations, can be expressed as:
-$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{B_{tot}}]\cdot(-\hat{y}),$ (2)
+$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{\tilde{B}_{tot}}]\cdot(-\hat{y}),$ (2)
 where $\mathcal{R}[\phi,\mathbf{v}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{v}$.
-For the applied rotating field in the xy-plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$.
-Therefore, $\mathbf{B_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$.
+In the co-rotating reference frame, the effective field after applying the rotating wave approximation is given by $\mathbf{\tilde{B}_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$.
 Here, $\hat{x}$, $\hat{y}$ and $\hat{z}$ are the unit vectors along the x, y and z axes, respectively.
-We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{B_{tot}}|$.
+We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{\tilde{B}_{tot}}|$.
 Eventually, the spin projections can be written as:
@@
 $P_{z}(t)=$ $\hat{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (4)
-where $\omega_{0}=\gamma|\mathbf{B_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
+where $\omega_{0}=\gamma|\mathbf{\tilde{B}_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
 $P_{x}(t)\approx\cos\left(\omega_{0}t+\gamma\frac{B_{m}b_{y}\cos\omega_{m}t-B_{m}b_{x}\sin\omega_{m}t}{\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}}\right)$ (5)
@@
 $\cos\omega_{m}t\cdot\frac{\gamma B_{z}-\omega_{m}}{\omega_{0}},$
-where $|\mathbf{B_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$.
+where $|\mathbf{\tilde{B}_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$.
 The residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift.
-The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 1 details the sign dependence).
+The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 1 details the sign dependence for $b_{x}$ and $b_{y}$ across various starting phases).
 Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
 The intuitive classical model underlying the operating principle of the FRF vector magnetometer is illustrated in Fig.
-1a.
+1.
 The scalar magnitude of the total magnetic field, $|\mathbf{B_{tot}}|$, depends on the angle between the rotating field $\mathbf{B_{m}}$ and the residual magnetic field $\mathbf{B_{0}}$.
@@
 The phases and amplitudes of these oscillations are influenced by the angle between $\mathbf{B_{0}}$ and $\mathbf{B_{m}}$.
-Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an in-phase oscillation, while a residual field in the y-direction induces an out-of-phase component.
-This model provides a fundamental description of the magnetometer’s operation.
-However, for precise measurements, several systematic effects—some even resulting from hyperfine interactions—require in-depth analysis to account for and minimize potential measurement biases.
-a) Experimental setup
-c) Phase shift analysis
-b) Four-shot rotating field scheme
-Figure 1: (a) Three coils determine the x, y, and z directions.
+Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an ou-of-phase oscillation, while a residual field in the y-direction induces an in-phase component.
+Figure 1: Experimental setup: Three coils determine the x, y, and z directions.
 The sensor head sits on a rotating stage with the cell positioned at the stage’s center.
 The sensor head can freely rotate in the xz and yz-planes.
-The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$.
-$\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field.
+$\alpha$ is the angle by which the probe beam moves away from the position where the probe beam is perpendicular to the static magnetic field in xz-plane.
+$\beta$ is the angle by which the pump beam moves away from the position where the pump beam is perpendicular to the static magnetic field in yz-plane.
 $\theta$ represents angle between the axis of rotation and the rotating field vector.
 Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode.
-(b) Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms.
+Figure 2: Four-shot rotating field scheme: Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms.
 When low after a falling edge, the pump laser and AC heater are turned off, and the probe beam laser measures the FID signal.
-There are four shots with different start phases of $B_{x}$ and $B_{y}$.The start phases of the rotating field play a significant role; therefore, we indicate them with vertical dashed lines.
+There are four shots with different start phases of $B_{x}$ and $B_{y}$.
+The start phases of the rotating field play a significant role as shown in Table 1; therefore, we indicate them with vertical dashed lines.
 This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA.
@@
 The signals during the preparation time are blanked out.
-(c) Upper: When a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in b) are analyzed by the MDA as a function of time.
+Figure 3: Phase shift analysis: Upper: when a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in Fig.
+2 are analyzed by the MDA as a function of time.
 Lower: when the magnetic fields in the x and y directions are roughly compensated.
 The first harmonic signals are suppressed and the second harmonic signal can be observed.
-II.2 Systematics
+This model provides a fundamental description of the magnetometer’s operation.
+However, for precise measurements, several systematic effects—some even resulting from hyperfine interactions—require in-depth analysis to account for and minimize potential measurement biases.
+2.2 Systematics
 To achieve high accuracy, systematic errors must be carefully studied and eliminated.
@@
 3 into a power series and examining the phase shifts to second order in $B_{m}$, we can identify two phase-shift contributions from $B_{m}$:
-[Berry’s phase shift] The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35].
+2.2.1 Berry’s phase shift
+The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35].
 As the magnetic field rotates slowly, the spins adiabatically follow the field.
@@
 where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig.
-1a.
+1.
 The Berry’s phase acquired over one complete period, given by Eq.
@@
 [35].
-[Second harmonic phase shift] Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
+2.2.2 Second harmonic phase shift
+Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
 $\phi_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (7)
@@
 Any asymmetry in the rotating field along the x and y-directions also produces a second harmonic signal instead of the first harmonic at $\omega_{m}$, ensuring that it does not affect the measurement of transverse magnetic fields.
-[Heading errors] The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36].
+2.2.3 Heading errors
+The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36].
 The pump beam heading error is well-studied and is primarily caused by nonlinear Zeeman splitting and the difference between Zeeman resonances in the two hyperfine ground states.
@@
 (More details in Sec.
-II.3).
-To clarify the elimination of the systematic dynamic heading error, we separate the dynamic heading error, $B_{DH}$, into two components: a static part ($B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{PD}$).
+2.3).
+To clarify the elimination of the systematic dynamic heading error, we separate the conventional dynamic heading error, $B_{D\!H}$, into two components: a static part ($B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{P\!D}$).
 Notably, the static heading error is independent of the rotating field phases.
@@
 Regarding the probe beam heading error, we define the angle that the probe beam rotates away from the x-axis as $\alpha$, with the sensor rotating in the xz-plane., as shown in Fig.
-1(b).
+1.
 The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$.
@@
 Converting it into a magnetic field unit, it corresponds to a fictitious magnetic field $-\omega_{m}/\gamma\cdot\tan\alpha$ along the x-direction, which we refer to as a probe beam heading error.
-[Eddy current] Another systematic error is induced by eddy currents.
+2.2.4 Eddy current
+Another systematic error is induced by eddy currents.
 The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field.
@@
 (More details in Sec.
-II.3)
-[Sign dependence of systematics] We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged.
+2.3)
+2.2.5 Sign dependence of systematics
+We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged.
 This characteristic offers an opportunity to eliminate these systematics.
@@
 We introduce a four-shot scheme (shown in Fig.
-1b) by altering the phases of the rotating field in the order listed in Table 1.
+2) by altering the phases of the rotating field in the order listed in Table 1.
 By averaging these four shots, the systematic errors can be canceled out.
@@
 4 $23\pi/$ 2 $2\pi$ + + - + - +
-II.3 Experimental results
+2.3 Experimental results
 To study systematics, a single-pass alkali vapor cell is placed at the center of a magnetic shield, which is a cubic cell with internal dimensions of 5$\times$5$\times$5 mm3.
@@
 As shown in Fig.
-1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction.
+1, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction.
 The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$.
@@
 The phase shift measuremment of one block (four shots) of data is shown in Fig.
-1c.
-The fitting model further fits the measured phase shifts $\delta\Phi(t)$ from the MDA as a function of measurement time to get: offset, slope ($\Phi_{s}$, Table A1 in the supplemental material details the components of the slope), amplitudes of the first harmonics ($b_{x}\propto\Phi_{x}$ and $b_{y}\propto\Phi_{y}$), second harmonics of $\omega_{m}$, and amplitudes of the first harmonics of the hyperfine frequency $\omega_{\text{hp}}$ (the difference in precession frequency for the F=1 and F=2 atoms induces an exponentially decaying sinusoidal phase shift at a frequency of $\omega_{\text{hp}}=2\pi\times 1.39$ kHz, see supplemental material Sec.
-I).
+3.
+The fitting model further refines the measured phase shifts, $\delta\Phi(t)$, for each shot from the MDA, modeling them as a function of measurement time to obtain: offset, slope ($\Phi_{s}$, Supplementary Table 2 in the Supplemental Information details the components of the slope), amplitudes of the first harmonics ($b_{x}\propto\Phi_{x}$ and $b_{y}\propto\Phi_{y}$), second harmonics of $\omega_{m}$, and amplitudes of the first harmonics of the hyperfine frequency $\omega_{hp}$ (the difference in precession frequency for the F=1 and F=2 atoms induces an exponentially decaying sinusoidal phase shift at a frequency of $\omega_{hp}=2\pi\times 1.39$ kHz, see Supplementary Note 1 in the Supplemental Information).
 $\begin{split}\delta\Phi(t)=&\Phi_{o\!f\!f}+\Phi_{s}\cdot t+\Phi_{x}\sin{\omega_{m}t}+\Phi_{y}\cos{\omega_{m}t}\\ &+\Phi_{2s}\sin{2\omega_{m}t}+\Phi_{2c}\cos{2\omega_{m}t}\\ &+\Phi_{hps}\sin{\omega_{hp}t}+\Phi_{hpc}\cos{\omega_{hp}t}\end{split}$ (10)
 […]
-II.3.1 Results of the systematics
+2.3.1 Results of the systematics
 The experimental results for the systematics are presented in Table 2.
+The Berry phase will result in a phase shift that adds to the total magnetic field measurement.
+When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq.
+6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table.
+2.
 Table 2: The measurement results and predictions are based on rotating matrix simulations and analytical models.
@@
 Static heading error $B_{S\!H}$ ($\beta\approx 24^{\circ}$) 2.7 nT 2.9 nT
-$|B_{PD}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$) 2.6 nT 2.8 nT
-[Berry’s phase] The Berry phase will result in a phase shift that adds to the total magnetic field measurement.
-When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq.
-6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table.
-2.
-[Second harmonic phase shift] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
-[Heading Errors] The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$.
+$|B_{P\!D}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$, $B_{m}$ initiates from Y) 2.6 nT 2.8 nT
+The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
+The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$.
 From Eq.
@@
 The slightly lower experimental value is attributed to incomplete spin polarization.
-(a) Probe beam heading error
-(b) Dynamic heading error
-(c) Eddy current due to different switches
-(d) Eddy current time constant
-Figure 2: (a) The measurement of the probe beam heading error.
-The solid points are experimental results, and the dashed lines are the theoretic plots based on our proposed model.
-(b) The magnetic field in the transverse direction is caused by the eddy current.
-The magnetic fields measured by both the Zero-Altering and Peak-Altering are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches.
-(d) The measurement of the eddy current time constant.
+The predicted value for $|B_{P\!D}|$ is based on approximate analytical model detailed in Supplementary Table 3
+Figure 4: Plots of systematic measurements: (a) Dynamic heading error: The spin precession plane adiabatically follows $\mathbf{B_{tot}}$.
+(b) Measurement of the probe beam heading error: The solid points represent experimental results, while the dashed lines correspond to theoretical predictions based on the proposed model.
+(c) Eddy currents generated by different rotating field alterations: Measurement of the transverse magnetic field (y-axis) induced by eddy currents due to zero-altering and peak-altering modulations.
+(d)Eddy Current Time Constant: Measurement of the eddy current time constant during a single panorama measurement of MDA.
 A rotating magnetic field can induce a rotating-field-phase-dependent heading error, $B_{P\!D}$.
 This effect is characterized in Table 3 with both experimental results and density matrix simulations [32].
-The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Secs.
-VIII and IX of the supplemental material.
-When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$.
-This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency.
-Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value.
-We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig.
-2(b).
-For example, when $\mathbf{B_{m}}$ begins along the x-axis, the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane.
-Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same.
-In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{PD}=0$.
-Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant.
-Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{DH}=0$.
-However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$.
-This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta\approx 2.9$ nT, and $B_{DH}=\mathcal{B_{H}}\cos\beta\sin\theta\approx 2.5~{}\mathrm{nT}$.
-Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{SH}$ remaining the same, but $B_{DH}=-\mathcal{B_{H}}\cos\beta\sin\theta\approx-2.5~{}\mathrm{nT}$.
-Table 1 provides a full overview of the sign dependence of the dynamic heading error.
-Table 3: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$.
-Here the static part of dynamic heading error $B_{SH}=0$.
+The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Supplementary Note 8 and Supplementary Tables of the supplemental Information.
+Table 3: Measured dynamic heading error, primarily influenced by $B_{PD}$, along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$.
+Here the static part of dynamic heading error $B_{SH}\approx 0$.
 “Exp” is the experimental data, “DM” is the density matrix simulation result.
-$\mathbf{B_{m}}$ initiates from Y-axis $\mathbf{B_{m}}$ initiates from X-axis ($\mathbf{B_{tot}}\perp\mathbf{S}$)
-$B_{tot}$ Shot No. Exp. (nT) DM (nT) Exp. (nT) DM (nT)
+$\mathbf{B_{m}}$ initiates from Y $\mathbf{B_{m}}$ initiates from X ($\mathbf{B_{tot}}\perp\mathbf{S}$)
+$B_{tot}$
+Shot
+No.
+Exp
+DM
+Exp
+DM
 1 2.6 3.0 0.1 0.0
@@
 Avg 0.1 0.1 0.1 0.0
+When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$.
+This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency.
+Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value.
+We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig.
+4a.
+For example, when $\mathbf{B_{m}}$ begins along the x-axis ($\phi_{x}=\pi/2$, $\phi_{y}=0$), the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane.
+Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same.
+In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{P\!D}=0$.
+Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant.
+Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{D\!H}=0$.
+However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$.
+This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$, and $B_{P\!D}=\mathcal{B_{H}}\cos\beta\sin\theta$.
+Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{S\!H}$ remaining the same, but $B_{P\!D}=-\mathcal{B_{H}}\cos\beta\sin\theta$.
+Table 1 provides a full overview of the sign dependence of the dynamic heading error.
 The experimental results for the dynamic heading error are smaller than the analytical model and density matrix simulation because the actual spin polarization is smaller than the simulation, which assumes full polarization.
-For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec.
-IX of the supplemental material.
+For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Supplementary Tables of the Supplemental Information.
 We further evaluated the dynamic heading error effect in the transverse directions.
@@
 For example, if the sensor is misaligned with an angle of $\alpha=1^{\circ}$, this results in an approximate 1.2 nT probe beam heading error in the x-direction.
-Additional data in Sec.
-IX of the supplemental material indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error.
+Additional data in Supplementary Tables of the Supplemental Information indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error.
 These systematic errors, however, can be eliminated through 4-shot averaging.
@@
 This error can introduce a systematic effect in transverse magnetic field measurements that depends on the angle $\alpha$, with its sign determined by the rotation direction of the rotating field.
-The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.2(a), with the theoretical basis provided in Eq.9.
+The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.
+4b, with the theoretical basis provided in Eq.
+9.
 “Theory 1” corresponds to anti-clockwise rotations (Shots 1 and 3 in Table 1), given by $-\omega_{m}/\gamma\cdot\tan\alpha$, while “Theory 2” represents clockwise rotations (Shots 2 and 4 in Table 1), given by $\omega_{m}/\gamma\cdot\tan\alpha$.
 The probe beam heading error can be effectively canceled by averaging the measurements from two opposite rotating field directions.
-[Eddy current] The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield.
-In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig.1b.
+The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield.
+In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig.
+2.
 We found that the way the rotating magnetic field is altered affects the eddy current and its induced magnetic field.
-When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig.1b, top left).
+When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig.
+2, top left).
 In this case, the average magnetic field caused by the eddy current equals zero.
 However, if the direction change occurs when the rotating magnetic field is at zero, termed “Zero-Altering” (Fig.
-1b, top right), it leads to a significant magnetic field caused by the eddy current.
-In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig.1b.
+2, top right), it leads to a significant magnetic field caused by the eddy current.
+In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig.
+2.
 In the “Zero-Altering” case, the second and fourth shots experience the maximum eddy current, while the first and third shots have smaller eddy currents, as they capture the decaying eddy current generated during the fourth and second shots.
 The sign of the eddy current-induced magnetic field depends on the rotation direction of the rotating field.
-A panorama measurement (80 shots) result from the MDA, shown in Fig.2(c), indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT.
+A panorama measurement (80 shots) result from the MDA, shown in Fig.
+4c, indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT.
 In contrast, Peak-Altering effectively suppresses the systematic effect of the eddy current.
@@
 In Fig.
-2(d), 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA.
+4d, 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA.
 Each shot measurement lasts 1/120 s.
@@
 To mitigate the eddy current’s systematic effect, we designed a waveform shown in Fig.
-1b to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering).
+2 to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering).
 By employing Peak-Altering and 4-shot averaging, we effectively reduce the eddy current’s systematic effect.
-II.3.2 Results of magnetic field sensitivities
-A magnetic field gradiometer measures the variation in magnetic field strength over a short distance, rather than the absolute field strength at a single location.
-This method is effective for detecting local magnetic changes while reducing the influence of uniform, distant magnetic interference [39].To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output.
-The laser, operating in pulsed mode, polarizes the spins along the z-axis to enhance spin alignment and is then turned off during the readout phase.
-Following this pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis.
+2.3.2 Results of magnetic field sensitivities
+A magnetic field gradiometer measures variations in magnetic field strength over a short distance, rather than the absolute field strength at a single location.
+This approach is highly effective for detecting local magnetic changes while minimizing the influence of uniform, distant magnetic interference [39].
+To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output.
+The laser operates in pulsed mode to polarize the spins along the z-axis, enhancing spin alignment, and is turned off during the readout phase.
+After the pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis.
+Notably, for synchronized optical pumping that polarizes spins approximately perpendicular to the external magnetic field, the average power requirement is typically only in the tens of milliwatts, while achieving comparable performance.
 Data acquisition is performed with two Carmel NK732 cards in place of the MDA, recording the zero-crossing data—including both frequency and phase information—from two magnetometers over a one-minute interval.
@@
 The gradiometer sensitivities are shown in Fig.
-3, with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ for the total field and 280 $\mathrm{fT/\sqrt{Hz}}$ for the $B_{x}$ and $B_{y}$ components, respectively.
+5, with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ for the total field and 280 $\mathrm{fT/\sqrt{Hz}}$ for the $B_{x}$ and $B_{y}$ components, respectively.
 Their noise spectrums are flat down to 1 Hz and 0.1 Hz, respectively.
 For comparison, a scalar sensitivity of 28 $\mathrm{fT/\sqrt{Hz}}$ was achieved without the application of a rotating field.
-Figure 3: Magnetic gradiometer sensitivity: The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields.
-The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer.
-The inset shows the frequency response for $B_{x}$ and $B_{y}$.
-III Discussion
+Figure 5: Magnetic gradiometer and polar angle sensitivities: The solid lines illustrate the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields.
+The dashed line depicts the sensitivity of the magnetometer in scalar mode.
+The inset highlights the frequency response for $b_{x}$ and $b_{y}$.
+The polar angle sensitivities, applicable only to $b_{x}$ and $b_{y}$, are shown on the right-side y-axis and are calculated by dividing $b_{x}$ and $b_{y}$ by the total static magnetic field, $B_{0}$.
+3 Discussion
 In summary, we present a FRF vector geomagnetic magnetometer.
@@
 The FRF vector magnetometer can be further improved to reach quantum-limited sensitivity [40].
-IV Methods
-IV.1 Shot noise limit
+4 Methods
+4.1 Shot noise limit
 To derive the shot noise limit for the FRF magnetometer, we utilize the CRLB [41] along with photon shot noise considerations specific to balanced polarimeter detection [42].
-Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Sec.
-IV of the supplemental material).
+Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Supplementary Note 4 of the Supplemental Information).
 For total magnetic field measurements, the sensitivity is expressed as
@@
 This noise limit also applies to scalar atomic magnetometers when analyzing the free induction decay (FID) signal.
-If the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
+When the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
 $\delta B_{tran}\geq\frac{2\sqrt{2}\omega_{m}\csc\theta}{\gamma k\sqrt{\Phi_{pr}}}.$ (12)
@@
 However, this assumption becomes inaccurate if the modulation frequency is not sufficiently high.
-Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Sections III and IV of the supplemental material.
+Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Supplementary Notes 4 and 5 of the Supplemental Information.
 The results show that when modulation frequency is sufficiently fast ($\omega_{m}\gtrsim\pi/T_{2}$), independent fitting yields comparable sensitivities.
@@
 Despite modifications to reduce current noise, it remains a primary noise source, limiting overall sensitivity.
-To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Sec. VI of the supplemental material for details):
+To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Supplementary Note 6 of the Supplementary Information for details):
 $\delta B_{tran\_S\!M}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}.$ (13)
@@
 Thus, the FRF vector magnetometer achieves transverse field sensitivity comparable to that of sequential modulation vector magnetometers while offering a higher bandwidth and reduced susceptibility to systematic effects.
+Data availability
+All data necessary to evaluate the conclusions of this study are included in the article, with supplementary data provided in tables and additional information included in the Supplementary Information.
+Any further details related to this study can be requested from the corresponding authors.
 Acknowledgements
-The authors thank Dmitry Budker, Conlon Lorcan Oneill for the valuable comments.
-The authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA), A*STAR Career Development Fund ( [^2] D800028), Italy-Singapore science and technology collaboration grant (R23I0IR042), Delta-Q (C230917004, Quantum Sensing), and Competitive Research Programme (NRF-CRP30-2023-0002).
-Author Contributions
+The authors thank Dmitry Budker and Conlon Lorcan Oneill for their valuable comments.
+All authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA) under contract No.
+HR001120C009 (M.R., T.K.).
 T.W.
-prepared the setup and conducted experiments for the FRF vector magnetometers, proposed and implemented the modulation, modeled spin evolution using rotation and density matrices, derived fundamental limits, analyzed results, and prepared the manuscript.
-M.R.
-supervised the setup and experiments, contributed to spin evolution modeling with rotation matrices, and discussed modulation and dynamic heading error.
-W.L.
-assisted with modeling static heading error and analyzing atomic polarization.
-M.L., T.K., and E.F.
-provided scalar magnetometer sensors, offered valuable experimental advice, and provided constructive comments for manuscript revisions.
-All authors reviewed the manuscript.
-Competing interests
-The authors declare no competing interests.
+acknowledges support from A*STAR Career Development Fund ( [^2] D800028)(T.W.), Italy-Singapore science and technology collaboration grant (R23I0IR042)(T.W.), Delta-Q (C230917004, Quantum Sensing)(T.W.), and Competitive Research Programme (NRF-CRP30-2023-0002)(T.W.).